### v1

# ToolACE-R: Tool Learning with Adaptive Self-Refinement

Abstract
Tool learning, which allows Large Language Models (LLMs) to leverage external tools for solving complex user tasks, has emerged as a promising avenue for extending model capabilities. However, current approaches primarily focus on data synthesis for fine-tuning LLMs to invoke tools effectively, largely ignoring how to fully stimulate the potential of the model. In this paper, we propose ToolACE-R, a novel method that introduces adaptive self-refinement for tool invocations. Our approach features a model-aware iterative training procedure that progressively incorporates more training samples based on the model’s evolving capabilities. Additionally, it allows LLMs to iteratively refine their tool calls, optimizing performance without requiring external feedback. To further enhance computational efficiency, we integrate an adaptive mechanism when scaling the inference time, enabling the model to autonomously determine when to stop the refinement process. We conduct extensive experiments across several benchmark datasets, showing that ToolACE-R achieves competitive performance compared to advanced API-based models, even without any refinement. Furthermore, its performance can be further improved efficiently through adaptive self-refinement. Our results demonstrate the effectiveness of the proposed method, which is compatible with base models of various sizes, offering a promising direction for more efficient tool learning.
ToolACE-R: Tool Learning with Adaptive Self-Refinement
Xingshan Zeng1, Weiwen Liu1, Xu Huang2, Zezhong Wang3, Lingzhi Wang4, Liangyou Li1, Yasheng Wang1, Lifeng Shang1, Xin Jiang1, Ruiming Tang1, Qun Liu1 1Huawei Noah’s Ark Lab 2University of Science and Technology of China 3The Chinese University of Hong Kong 4Harbin Institute of Technology, Shenzhen zeng.xingshan@huawei.com, liuweiwen8@huawei.com, xuhuangcs@mail.ustc.edu.cn
1 Introduction
Tool learning, which enables Large Language Models (LLMs) to leverage external tools to address complex user requirements, has gained increasing attention. With tool integration, LLMs can access up-to-date information, perform intricate computations, and utilize third-party services, significantly expanding their capabilities beyond simple natural language communication with humans Qu et al. (2024b). While tool invocation requires the LLMs to demonstrate strong understanding, reasoning, and instruction-following skills, customized fine-tuning is currently the dominant approach for enabling models to call external tools Qin et al. (2023); Patil et al. (2023); Liu et al. (2024b, c).
Due to the limited availability of high-quality data, existing research has primarily focused on developing effective and efficient methods for data synthesis using advanced models Liu et al. (2024b, c); Wang et al. (2024b); Abdelaziz et al. (2024). However, data synthesized by advanced models from different sources can lead to compatibility issues. Specifically, when synthesized samples exceed the model’s current knowledge, they may undermine the model’s performance or lead to hallucinations Ren et al. (2024); Kang et al. (2024). As a result, determining which training data samples are appropriate for a given model remains an unresolved challenge.
Additionally, the potential of maximizing the model’s own capabilities remains under exploration in the field of tool learning. It has been shown that a model’s reasoning performance can be improved by scaling test-time computation Brown et al. (2024); Snell et al. (2024). One promising approach involves iteratively refining the model’s output multiple times and then selecting the majority answer Snell et al. (2024). However, user queries vary in complexity, ranging from simple to highly complex. Scaling test-time computation for overly simple and straightforward questions is inefficient, as models can often answer them correctly without additional refinement. Therefore, an adaptive strategy for scaling test-time computation is needed to dynamically determine the appropriate level of effort required for each question.
To address the challenges outlined above, we propose ToolACE-R, a method that enables adaptive self-refinement of tool invocations through a customized iterative training procedure. Specifically, we introduce a model-aware iterative training approach that facilitates LLMs in learning self-refinement for tool invocation in an iterative fashion, enhanced by model-based data selection. Additionally, by constructing refining data samples that preserve identical cases, LLMs are more likely to learn appropriate stopping criteria during the iterative process. Building on this training procedure, we further integrate adaptive self-refinement into iterative inference for inference time scaling, allowing models to autonomously determine when to halt the iteration, thereby improving scaling efficiency.
We have conducted extensive experiments on several representative tool-calling benchmarks, such as the Berkeley Function Call Leaderboard (BFCL) Yan et al. (2024) and API-Bank Li et al. (2023), to assess the effectiveness and efficiency of ToolACE-R.The experimental results show that ToolACE-R achieves competitive performance compared to advanced API-based models like GPT-4o, and can be further enhanced efficiently through adaptive self-refinement. Additional analyses highlight the impact of the proposed modules and examine the generalizability of the method across different model backbones and sizes.
The key contributions of this work are summarized as follows:
  • We propose a model-aware iterative training procedure, enabling LLMs learning adaptive self-refinement for tool learning, which maximizes the model potential without any feedback.
  • We integrate adaptive self-refinement into iterative inference, enabling models to autonomously decide when to halt the refinement process and allowing for more efficient use of computational resources when scaling test-time compute.
  • We conduct extensive experiments on several representative tool calling benchmarks, revealing effectiveness and efficiency of our method.
2 Related Work
Tool Learning.
Integrating external tools can enhance LLMs’ capabilities, enabling more specialized, precise, and reliable problem-solving for users (Qin et al., 2023). Tool learning methods generally fall into two categories, tuning-free and tuning-based approaches. Tuning-free methods employ prompting techniques, such as in-context tool-use examples, without additional training Mialon et al. (2023); Hsieh et al. (2023); Ruan et al. (2023). A notable example method is ReAct Yao et al. (2023), which can alternate between reasoning and action for complex tasks. Recent work Qu et al. (2024a) improves tool understanding by iteratively refining documentation. Tuning-based methods, which directly enhance tool calling ability via specialized fine-tuning, have recently gained much more attention Qin et al. (2023); Schick et al. (2023); Patil et al. (2023); Tang et al. (2023); Abdelaziz et al. (2024). Due to data scarcity, a large amount of them focus on data synthesis for improved generation procedures Liu et al. (2024b, c); Wang et al. (2024b), often overlooking the model’s inherent potential. While some others integrate tool feedback for refinement Du et al. (2024); Wang et al. (2024a), our approach is more streamlined, directly refining outputs in an iterative manner without external or any textual feedback to maximize model efficiency.
Data Selection.
Research indicates that selecting high-quality training samples is essential for fine-tuning LLMs Albalak et al. (2024). When the base model is sufficiently powerful, a small set of high-quality data can effectively harness the model’s potential, rather than relying on large quantities of general data Zhou et al. (2023); Liu et al. (2023). While earlier works primarily emphasize general data quality aspects, such as diversity and complexity, recent studies advocate for model-specific data selection Du et al. (2023); Li et al. (2024). This kind of approach is based on the observation that data distributions that significantly deviate from the base model’s are challenging for the model to learn from and may even degrade the performance Ren et al. (2024). Our method aligns with this perspective, proposing the use of pass@k metric to identify and select suitable training samples based on the model to be trained.
Self Refinement.
Previous research has demonstrated that LLMs can refine their own outputs through either self-feedback Madaan et al. (2023); Weng et al. (2023) or external feedback Qu et al. (2024c); Xu et al. (2024). Nevertheless, it remains challenging for LLMs to assess the correctness of their refined output autonomously Huang et al. (2024), which means that scaling test-time compute relying on iterative refinement still often depend on post-processing techniques, such as self-consistency Snell et al. (2024). In this work, we propose a simple yet effective method that enables models to adaptively self-refine their outputs, thereby improving scaling efficiency.
Figure 1: The overview of ToolACE-R, which includes two components, for training and inference, respectively.
3 Methodology
3.1 Problem Formulation
Tool learning aims to address the following problem: given a user query $q$ and a set of candidate tools $T=\{t_{1},t_{2},\dots,t_{n}\}$, the objective is to generate the correct tool invocations. This includes selecting the most appropriate tools and extracting the suitable parameter values. Specifically, the goal is to produce:
$A=[t_{1}(a_{1}),\dots,t_{m}(a_{m})]=f_{\theta}(\langle q,T\rangle)$ (1)
where $t_{j}$ and $a_{j}$ represent the $j$-th invoked tool and its corresponding argument, respectively, with $1\leq j\leq m$ and $m$ being the total number of tool invocations needed. The function $f_{\theta}(\cdot)$ denotes the generation process of the LLM with parameters $\theta$.
The training sample for tool learning is typically formalized as a triplet consisting of a user query, candidate tools, and the ground-truth tool invocation: $\{\langle q,T\rangle,A\}$, where $q$ and $T$ serve as input and $A$ should be output.
Our ToolACE-R is trained with a model-aware iterative training pipeline and can iteratively self-refine during inference in an adaptive manner. Figure 1 shows the overall procedure.
3.2 Model-Aware Iterative Training
Our model training starts with an instructed base model, ensuring that it has basic instruction following capability for our model-aware module. For training, we first gather a set of off-the-shelf training samples. Instead of using them directly, we apply a selection criterion to curate the data, which is then combined with generated self-refinement data. This curated and refined data is integrated into an iterative training process.
Data Selection Criterion.
Inspired by self-improvement techniques in existing LLM reasoning research Zelikman et al. (2022), where LLMs are trained using their own outputs, we introduce a pass@k criterion to select the most suitable training samples for each base model – those the model can answer (almost) correctly within at most $k$ tries. This approach may seem counterintuitive, as one might expect models to focus on learning from samples they struggle with. However, recent research Ren et al. (2024) has highlighted that models are more likely to perform self-alignment rather than learning new information during the fine-tuning stage. Based on this insight, we adopt this pass@k metric for data selection.
Specifically, for each sample $\{\langle q,T\rangle,A\}$, the base model to be trained first generate the invocation up to $k$ times. If any of the responses approaches the groundtruth $A$, we retain it as an official training sample:
$R(\langle q,T\rangle,A)=\begin{cases}1&\text{if }\exists\,j\in\{1,2,..,k\}\,\text{s.t. }\\ &f_{\theta_{0}}(\langle q,T\rangle)_{j}\approx A,\\ 0&\text{otherwise}.\end{cases}$ (2)
Here $R(\langle q,T\rangle,A)=1$ indicates the training sample will be retained. $\theta_{0}$ represents the base model before training and $f_{\theta_{0}}(\cdot)_{j}$ means the $j$-th generation with the base model.
A key aspect is how to define when a response "approaches" $A$. In our preliminary experiments, we found that if we restrict them to be exactly equal, the resulting training set still has an effect but limits the potential for performance improvement (see Table 4). However, by loosening this constraint, more suitable training samples are retained and the model’s potential can be more effectively harnessed. Therefore, we use the concept of "approaching" in this context. For any two invocations $A$ and $A^{\prime}$, we define $A\approx A^{\prime}$ if the used tools are exactly the same and each invocation only contains at most 1 different parameter values.
Self-Refine Data Construction.
To establish the model’s self-refinement capability, we additionally construct self-refine data based on the training samples. Specifically, we create a self-refine sample as a multi-turn interaction, where the first turn is the initial response from the model, and the second turn includes a refinement prompt from the user followed by the model’s refined answer: $\{\langle q,T\rangle,A_{1},r,A_{2}\}$. We apply the same selection criterion as described above, retaining the self-refinement training sample only when $A_{2}\approx A$:
$R(\langle q,T\rangle,A_{1},r,A)=\begin{cases}1&\text{if }\exists\,j\in\{1,2,..,k\}\,\text{s.t. }\\ &f_{\theta_{0}}(\langle q,T\rangle,A_{1},r)_{j}\approx A,\\ 0&\text{otherwise}.\end{cases}$ (3)
Here, we replace $A_{2}$ with $A$ to create the refined training sample: $\{\langle q,T\rangle,A_{1},r,A\}$.
Notably, there may be samples where $A_{1}=A$, meaning the refined answer is identical to the previous answer. We retain these samples to enable the model to learn that "no further changes are needed when the answer is perfect." This is a crucial aspect of establishing the model’s adaptive, iterative self-refinement, and we will provide further details in Section 3.3.
Iterative Training.
Given the selected training samples (including self-refine samples), we apply supervised fine-tuning to the base model for one epoch. This constitutes one iteration. In subsequent iterations, we treat the trained model as the new base model, re-construct the self-refine samples, and re-select those that are suitable for fine-tuning the updated model. The model is then further trained using the newly selected samples.
Ideally, with each iteration, the model’s performance improves, leading to an increase in the number of selected samples. The iteration process continues until the number of selected samples no longer increases, indicating that the model has saturated on this set of training samples. The resulting model will then be our final ToolACE-R model.
3.3 Adaptive Self-Refine Iterative Inference
With model-aware iterative training, our ToolACE-R model can directly generate tool invocations based on a user query and candidate tools, and refine its answer when necessary.
We propose an adaptive self-refine procedure that enables ToolACE-R to iteratively refine its own output and autonomously determine when to halt the iteration. The details are introduced in Algorithm 1. Specifically, given a user query $q$ and candidate tools $T$, ToolACE-R first generates the tool invocation $A_{0}$. In each subsequent iteration, ToolACE-R leverages the answer from the previous iteration as context and refines it. As mentioned earlier, ToolACE-R has learned that when the answer is sufficiently accurate, it can produce the same answer in subsequent refinement. Therefore, the iteration continues until the answers from two consecutive iterations are identical. We refer to this procedure as adaptive self-refine, where the model itself determines when to stop the iterations. To prevent infinite iteration, we set a maximum iteration limit $n$.
Algorithm 1 Adaptive Self-Refine
0:  Model $\theta$, query $q$, tools $T$, refine prompt $r$, time $n$
0:  Refined final answer $A$
1:  $A_{0}=f_{\theta}(\langle q,T\rangle)$  // First get the direct answer $A_{0}$
2:  for $i=1$ to $n$ do
3:     $A_{i}=f_{\theta}(\langle q,T\rangle,A_{i-1},r)$  // Get the refined answer
4:     if $A_{i}==A_{i-1}$ then
5:        return  $A_{i}$  // Adative stopping
6:     end if
7:  end for
8:  return  $A_{n}$  // $n$ is the maximum iteration time
Through adaptive self-refine, ToolACE-R is able to improve its performance, particularly when facing difficult cases.
4 Experiments
Non-Live Live Overall
Models Simple Multiple Parallel
Parallel
Multiple
Simple Multiple Parallel
Parallel
Multiple
Non-live Live Overall
GPT-4o-2024-11-20 79.42 95.50 94.00 83.50 84.88 79.77 87.50 75.00 88.10 80.75 84.43
Gemini-1.5-Pro-002 78.33 93.50 92.5 90.00 82.95 77.40 87.50 83.33 88.58 78.68 83.63
GPT-4o-mini-2024-07-18 80.08 90.50 89.50 87.00 81.40 76.73 93.75 79.17 86.77 77.87 82.32
GPT-3.5-Turbo-0125 77.92 93.50 67.00 53.00 80.62 78.63 75.00 58.33 72.85 78.61 75.73
DeepSeek-V3 78.67 95.5 91.00 91.50 83.72 82.15 81.25 62.50 89.17 82.09 85.63
Llama3.1-70B-Inst 77.92 96.00 94.50 91.50 78.29 76.16 87.50 66.67 89.98 76.53 83.26
Llama3.1-8B-Inst 71.50 93.50 86.50 86.00 73.26 68.95 56.25 50.00 84.37 69.28 76.83
Qwen2.5-7B-Inst 70.50 94.00 90.50 82.50 74.81 71.89 62.50 62.50 84.38 72.17 78.28
ToolACE-8B (FC) 76.67 93.50 90.50 89.50 73.26 76.73 81.25 70.83 87.54 76.02 81.78
Hammer2.1-7B (FC) 78.08 95.00 93.50 88.00 76.74 77.4 81.25 70.83 88.65 77.20 82.92
xLAM-7B-r (FC) 74.25 95.50 81.00 73.50 72.09 74.93 50.00 62.50 81.06 73.87 77.47
ToolACE-R (FC) 79.92 96.50 93.50 93.00 81.78 81.01 87.50 83.33 90.73 81.27 86.00
  + Adaptive Self-Refine 81.17 96.00 94.00 93.00 82.95 81.58 87.50 83.33 91.04 81.94 86.49
Table 1: Accuracy comparison on BFCL (Last updated on 2025-01-18). The best results in each category are marked in bold. The second best results are underlined. FC indicates the models are fine-tuned for function calling.
4.1 Experimental Settings
Datasets and Models.
We leverage a subset of ToolACE Liu et al. (2024b) training data as our original full training dataset. The dataset contains multiple types of tool calling data, including cases with multiple candidate tools and parallel function callings, in both Chinese and English. For base model to be trained, we select LLaMA3.1-8B-Instruct AI@Meta (2024) for our main experiments. We also conduct experiments on other base models, including Qwen2.5-Instruct-series Yang et al. (2024) (0.5B, 1.5B, 3B and 7B) and Mistral-7B-Instruct-v0.2 Jiang et al. (2023), to validate the generalizability of ToolACE-R. We compare with the state-of-the-art API-based models, including GPT-series [^1] https://chatgpt.com and Gemini-1.5-Pro [^2] https://gemini.google.com/, as well as open-source models like DeepSeek-V3 Liu et al. (2024a), Llama3.1-8B/70B-Instruct AI@Meta (2024) and Qwen2.5-7B-Instruct Yang et al. (2024). We also compare with fine-tuned tool calling models like ToolACE-8B Liu et al. (2024b), Hammer2.1-7B Lin et al. (2024) and xLAM-7b-r Zhang et al. (2024).
Benchmarks and Evaluation.
We conduct experiments on several representative benchmarks, including the Berkeley Function Call Leaderboard (BFCL) Yan et al. (2024), ACEBench Chen et al. (2025), API-Bank Li et al. (2023) and ToolAlpaca Tang et al. (2023). To fully explore the model’s potential while minimizing external interference and simplifying data preparation, we mainly evaluate performance on single-turn tool calling queries. For BFCL, we evaluate both the Non-live and Live subsets, which correspond to synthetic test cases and real-world scenarios, respectively. Each subset includes four categories: Simple, Multiple, Parallel, and Parallel Multiple. Simple and Multiple cases both involve a single invoked tool, but in Multiple cases, there are multiple candidate tools. Parallel and Parallel Multiple cases require invoking multiple tools simultaneously. For ACEBench, we evaluate only the English normal category, excluding multi-turn cases. This includes the test samples in atom category and single-turn category. For API-Bank and ToolAlpaca, we focus on the first step tool calling samples, disregarding further tool or retrieval feedback. Performance is reported in terms of accuracy for all benchmarks.
Implementation Details.
Given resource constraints, we employ the parameter-efficient fine-tuning method LoRA Hu et al. (2022). All model modules are configured for LoRA fine-tuning, with a rank of 16 and an alpha value of 32. Training is performed with a global batch size of 64 and a learning rate of $1\times 10^{-4}$, following a cosine learning rate schedule with a warmup ratio of 0.1. For data selection, we set the generation temperature to 1.0 to encourage diversity and allow the model to explore a broader range of outputs. Each sample is generated up to 8 times (pass@8) to ensure potential. For the constructed refined data, we apply a loss mask to the first assistant turn of the conversation, as it may involve an incorrect tool invocation, making it unsuitable for the model to learn from this turn. During evaluation, we adopt greedy search to ensure stability and impose a maximum iteration limit of $n=5$.
4.2 Main Results
Results on BFCL.
Table 1 presents the comparison results on BFCL, including detailed results for each category and the overall average performance. The following key observations can be made:
  • The performance gap between API-based and open-source models is minimal. Large open-source models like DeepSeek-V3 and Llama3.1-70B-Inst show comparable or even superior overall performance to GPT-4o and Gemini-1.5-Pro.
  • Specialized fine-tuned models significantly benefit from domain-specific fine-tuning, allowing smaller models (7/8B) to compete with larger general models like Llama3.1-70B-Inst.
  • Our model, ToolACE-R which leverages our proposed iterative training procedure, achieves the best overall performance, surpassing both large API-based and open-source models.
  • The adaptive self-refinement mechanism further enhances the performance of ToolACE-R, with larger improvement observed in more complex categories (Live), highlighting the effectiveness of our inference method.
Models ACEBench APIBank ToolAlpaca
GPT-4o-2024-11-20 87.00 77.16 83.87
GPT-4o-mini-2024-07-18 81.63 73.10 83.87
Llama3.1-8B-Inst 50.63 60.41 79.03
ToolACE-R 83.88 75.89 85.48
  + Adaptive Sef-Refine 84.00 81.22 88.71
Table 2: Accuracy comparison on ACEBench, API-Bank and ToolAlpaca benchmarks.
Results on More Benchmarks.
To provide a more comprehensive evaluation, we continue to conduct experiments on three other representative benchmarks ACEBench, API-Bank, and ToolAlpaca, focusing on single-turn tool-calling scenarios. The results are summarized in Table 2.
On ACEBench, ToolACE-R demonstrates an impressive absolute improvement of over 30% compared to the base model, Llama3.1-8B-Inst, highlighting the effectiveness of our training pipeline. However, the benefit of Adaptive Self-Refine is marginal, likely due to the presence of particularly challenging cases where the model struggles to correctly refine its answers.
For both API-Bank and ToolAlpaca, ToolACE-R achieves substantial improvements. Moreover, when combined with Adaptive Self-Refine, the performance sees further gains, even surpassing GPT-4o. These results validate the potential of Adaptive Self-Refine to progressively refine tool invocation answers based solely on the model’s own capabilities, opening up new possibilities for scaling inference in tool learning scenarios.
Models Non-live Live Overall
Llama3.1-8B-Inst (Base) 84.37 69.28 76.83
ToolACE-R (Full) 91.04 81.94 86.49
  - Adaptive Self-Refine 90.73 81.27 86.00
    - Iterative Training 88.12 80.24 84.18
      - Data Selection 88.96 81.13 85.04
        - Refine Data 88.75 80.09 84.42
Table 3: Ablation Study of our proposed modules.
Ablation Study.
To investigate the impact of the proposed modules, we conduct an ablation study by progressively removing them. The corresponding evaluation results on the BFCL dataset are shown in Table 3. As observed, removing both adaptive self-refinement and iterative training leads to a noticeable performance degradation. However, further removing data selection does not result in a continued drop in performance. In fact, performance improves. This suggests that model-aware data selection should be combined with iterative training to fully leverage the entire training set, as the initial data selection (performed by the base model) typically only captures a small subset of the available data (mostly easy ones). Iterative selection is needed to capture more complex cases, as the trained model is improved. Finally, removing the constructed refine data also negatively impacts performance, validating that self-refine data acts as a form of data augmentation, also enhancing the model’s direct output.
4.3 Effects of Iterative
In this subsection, we examine the effects of our two iterative modules, i.e., model-aware iterative training and adaptive self-refine iterative inference.
Figure 2: Performance of each iteration during training. “Amount” indicates the amount of training samples in that iteration with our data selection module. Iteration 0 means the model before training, i.e., Llama3.1-8B-Inst.
Iterative Training.
Figure 2 illustrates the results of the iterative training process, showing both the accuracy performance and the amount of training data selected at each iteration. Iteration 0 corresponds to the base model, Llama3.1-8B-Inst. As depicted, the quantity of selected samples increases monotonically across iterations, suggesting that the model’s performance on the training set improves, as the selected samples are those the model is more likely to answer correctly (as indicated by pass@k). On the other hand, accuracy improves rapidly during the first two iterations, reaching its peak at iteration 3. Subsequently, accuracy plateaus, with no further improvement in later iterations. This suggests that excessive training could lead to overfitting on the training set.
Figure 3: Performance of different iterative self-refine methods. “Vanilla Self-Refine” always picks the final answer at each iteration. “Adaptive Self-Refine” is our proposed method, while “Self-Consistency” selects the majority answer during the iteration.
Iterative Inference.
To assess the effects of our adaptive self-refine method for iterative inference, we conducted additional experiments comparing it with two other inference-time scaling approaches. The first, termed "Vanilla Self-Refine," always selects the last answer at each iteration as the final answer. The second method, "Self-Consistency at Itr n", chooses the majority answer from the answers collected during iterations 1 to n. The results of these inference methods on BFCL, all based on the same model, are shown in Figure 3.
As observed, the performance of "Vanilla Self-Refine" fluctuates across iterations, indicating that the model does not always refine the answer correctly. In contrast, our "Adaptive Self-Refine" method demonstrates more stability across iterations, suggesting that adjacent identical answers serve as a reliable signal for stopping further iterations. The "Self-Consistency at Itr n" method likely represents an upper bound for answer selection, with our method achieving competitive performance compared to this strong baseline. Furthermore, the performance of our method remains consistent across different iteration counts (3, 5, 10), likely because most cases terminate early, well before reaching the maximum iteration limit. This observation is supported by the average iteration time of 2.4, as indicated by our statistics on BFCL.
4.4 Effects of Data Selection
In this subsection, we investigate the effects of our model-aware data selection module. First, we explore the impact of the $k$ value when using pass@k for sample selection. Next, we examine how the use of data selected by different models affects model performance, highlighting the impact of model-aware selection.
Models Direct Refine Amount
Select with Pass@8
  First Iteration 80.24 80.61 120K
  Best Iteration 81.27 81.94 150K
Select with Pass@1
  First Iteration 78.90 78.53 94K
  Best Iteration 79.57 80.01 143K
Table 4: Accuracy performance on BFCL Live category when data selection with Pass@8 and Pass@1. “Direct” and “Refine” indicate the results with direct output and adaptive self-refine, respectively. “Amount” corresponds to the amount of selected training samples.
Data Selection with Pass@k.
In our data selection criterion, we retain training samples that the base model can answer (approximately) correctly within at most $k$ tries (pass@k). In this section, we examine the effects of varying $k$ values. Specifically, we report the evaluation results using pass@8 and pass@1 in Table 4. As shown, although the final number of selected samples is similar between the two methods (150K vs. 143K), the accuracy performance for pass@8 is considerably higher. The performance gap is already pronounced in the first iteration, suggesting that a more lenient selection criterion plays a crucial role in maximizing model potential and shaping the model’s distribution. Notably, this effect does not diminish over additional iterations. Based on these findings, we select a larger $k$ value for our experiments.
Model-aware Selection.
In our data selection process, we propose selecting training samples in a model-aware manner, ensuring that a sample is chosen only when the base model can answer it (almost) correctly. This is based on the assumption that samples aligning with the model’s distribution are more suitable for training. To further assess the impact of such a model-aware approach, we conduct an additional experiment. Instead of using data selected by the base model Llama3.1-8B-Inst, we train Llama3.1-8B-Inst with data selected by Qwen2.5-3B-Inst and Qwen2.5-7B-Inst, as well as with the intersection of all three models’ selected data – i.e., only the samples that all three models perform well on are included in the dataset. The accuracy results of the trained models, along with the corresponding training data amounts, are presented in Table 5. As shown, performance drops when using data selected by other models, even when the data volume is larger. The intersection dataset, which contains less data, achieves comparable results. This observation underscores the benefits of model-aware data selection.
Models Accuracy Amount
ToolACE-R 80.61 120K
Data selected with Qwen2.5-3B 79.50 125K
Data selected with Qwen2.5-7B 80.38 134K
Intersection of Three Datasets 80.27 108K
Table 5: Accuracy performance on BFCL Live category in the first interation when using different selected data to train Llama3.1-8B-Inst, along with the data amount. “Qwen2.5-3/7B” is short for Qwen2.5-3/7B-Inst.
4.5 Further analysis
In this subsection, we examine the generalizability of our method by varying the base model and exploring the scaling effects across different model sizes.
Models Non-live Live Overall
Raw (Qwen2.5-7B-Inst) 84.38 72.17 78.28
  + Direct Data 88.29 78.39 83.34
  + ToolACE-R 89.23 81.20 85.21
Raw (Mistral-7B-Inst-v0.2) 55.25 55.07 55.16
  + Direct Data 84.06 70.17 77.12
  + ToolACE-R 87.98 80.01 84.00
Table 6: Accuracy performance of Different backbones on BFCL. “+ Direct Data” indicates the model directly trained with the full original set, while “+ ToolACE-R” is the result with our method, both iterative training and adaptive self-refine are applied.
Different Backbones.
To assess the generalizability of ToolACE-R, we conduct experiments with different base models, including Qwen2.5-7B-Inst and Mistral-7B-Inst-v0.2. The results, along with the original performance and performance after training with the full original training set, are presented in Table 6. As shown, fine-tuning with the full training set leads to a significant performance boost for both models, particularly for the relatively weaker Mistral-7B-Inst-v0.2, which sees an improvement of over $20\%$. Moreover, our method further enhances performance, using exactly the same original data, demonstrating its effectiveness.
Model Size Scaling.
Figure 4: Performance of models in Qwen2.5-Series. “Base” refers to the Qwen2.5-Instruct models, while “Direct” and “Adaptive Self-Refine” represent the results of models trained using our method and evaluated with direct output and Adaptive Self-Refine, respectively.
We further examine whether our method can be applied to models of varying sizes and whether performance improves as model size increases. For this, we use Qwen2.5-Instruct models of different sizes, including 0.5B, 1.5B, 3B, and 7B, as base models. The results, shown in Figure 4, display both the original performance and the performance after applying ToolACE-R. The direct output of ToolACE-R consistently improves performance across all sizes, particularly for smaller models like 0.5B and 1.5B, and the Adaptive Self-Refine inference technique further enhances these results. This demonstrates the effectiveness of ToolACE-R across a range of model sizes.
5 Conclusion
In this paper, we introduce ToolACE-R, a tool learning technique designed to enable large language models (LLMs) to effectively leverage external tools. ToolACE-R consists of two key components: model-aware iterative training and adaptive self-refine iterative inference. Model-aware iterative training maximizes the model’s potential by iteratively selecting the most appropriate samples for training. Adaptive self-refine further enhances performance by allowing the model to autonomously determine when to stop refining its own answers during iterative inference. Experimental results demonstrate that ToolACE-R consistently improves performance across various tool-calling benchmarks and is effective with different backbone models. Additionally, our analyses highlight the impact of the proposed modules on model performance.
Limitations
We summarize the limitations in the following two points:
First, due to resource constraints, we conduct experiments only on models with up to 8B parameters, using LORA. As shown in the scaling performance results in Figure 4, we have not yet observed any bottlenecks preventing larger models from benefiting from our ToolACE-R method.
Second, our focus is on single-turn tool-calling scenarios, excluding the multi-turn and multi-step cases for simplicity. However, we believe that single-turn cases are a fundamental aspect of tool learning. The significant improvements observed across several representative benchmarks demonstrate the effectiveness of our method. We plan to explore ToolACE-R in more complex scenarios in future work.
Appendix A Appendix
A.1 Detailed Introduction to Benchmarks
Below we introduce the used benchmarks with more details:
  • BFCL Yan et al. (2024): BFCL has progressed over time and is currently at version 3, which includes three main categories: non-live, live, and multi-turn. This paper focuses on the non-live and live categories, which primarily consist of Python-style tool call data. These are further subdivided into four types: Simple, Multiple, Parallel, and Parallel Multiple.
    • Single: A single evaluation is the simplest and most frequently encountered format, where the user provides a single function document, triggering exactly one tool call.
    • Multiple: The multiple category involves a user query that selects a single tool call from 2 to 4 available function documents. The model must choose the most appropriate tool to call based on the user’s context.
    • Parallel: A parallel test sample involves triggering multiple tool calls simultaneously in response to a single user query. The model must determine how many calls are needed.
    • Parallel Multiple: The parallel multiple cases combine the concepts of both parallel and multiple. Here, the model is provided with several function documents, and each corresponding tool call may be invoked zero or more times.
  • ACEBench Chen et al. (2025): ACEBench is a comprehensive tool-use benchmark that offers more detailed granularity. It is categorized into three main types: Normal, Special, and Agent. In this work, we focus on the Atom and Single-turn sub-categories within the Normal category. Atom cases involve a set of APIs that contain specific parameter types, such as booleans, enumerations, numbers, lists, and objects. The Single-turn category includes both single and parallel cases.
  • API-Bank Li et al. (2023): API-Bank is a dialogue-style tool call dataset, consisting of two settings: Call and Retrieve + Call. In this dataset, the model is tasked with invoking predefined local Python tools based on the user’s requirements in the dialogue. Accuracy is assessed by comparing whether the tool’s returned values match the ground truth. In this work, we focus on the Call setting in our experiments.
  • ToolAlpaca Tang et al. (2023): It introduces a multi-agent simulator that uses GPT-4 to simulate the return values of tools. The model can make modifications and re-call the tool based on the returned values (e.g., error messages). Ultimately, GPT-4 is used to evaluate the accuracy. In our experiments, we does not leverage the returned values for refinement for consistency.
A.2 Detailed Results in ACEBench
Table 7 displays the detailed results in each sub-category of ACEBench. As can be seen, in all sub-categories, ToolACE-R performs much better than its base model Llama3.1-8B-Inst. The improvement of Adaptive Self-Refine is not significant.
Atom Single-Turn Overall
Models Bool Enum Number List
Object
Short
Object
Deep
Single Parallel Overall
GPT-4o-2024-11-20 94.00 94.00 98.00 94.00 64.00 96.00 84.00 72.00 87.00
GPT-4o-mini-2024-07-18 88.00 94.00 98.00 90.00 44.00 92.00 77.00 70.00 81.63
Llama3.1-8B-Inst 50.00 74.00 48.00 60.00 30.00 44.00 55.00 44.00 50.63
ToolACE-R 88.00 96.00 94.00 94.00 50.00 92.00 83.00 74.00 83.88
  + Adaptive Self-Refine 88.00 96.00 94.00 94.00 50.00 94.00 84.00 72.00 84.00
Table 7: Detailed accuracy comparison on ACEBench.
A.3 Prompts
We show the detailed system prompt and refine prompt in Figure 5 and Figure 6, respectively. Note that the displayed prompts are those used during inference. For the training data, we diversify the system prompts to ensure better generalizability.
A.4 Data example
Figure 7 is an example of our refine data. The model will only learn the last turn, with previous turns as context.
You are an expert in composing functions. You are given a question and a set of possible functions. Based on the question, you will need to make one or more function/tool calls to achieve the purpose. If none of the function can be used, point it out. If the given question lacks the parameters required by the function, also point it out. You should only return the tool call in tools call sections. If you decide to invoke any of the function(s), you MUST put it in the format of: [[[[func_name1(params_name1=params_value1, params_name2=params_value2,…), func_name2(params)]]]] You should not include any other text in the response. Here is a list of functions in JSON format that you can invoke: {candidate tools} {other information}
Figure 5: The system prompt for tool calling. “{candidate tools}” is replaced with actual tool descriptions. “{other information}” is replaced with some other information in the case, e.g., date. If no other information, it is null.
Please refine your answer. Directly output the refined answer, or the original answer if you think it is already perfect.
Figure 6: The refine prompt for tool calling.
[System] You are an expert in composing functions. You are given a question and a set of possible functions. Based on the question, you will need to make one or more function/tool calls to achieve the purpose. If none of the function can be used, point it out. If the given question lacks the parameters required by the function, also point it out. You should only return the tool call in tools call sections. If you decide to invoke any of the function(s), you MUST put it in the format of: [[[[func_name1(params_name1=params_value1, params_name2=params_value2,…), func_name2(params)]]]] You should not include any other text in the response. Here is a list of functions in JSON format that you can invoke:
    "name": "bookFlight",
    "description": "Book a flight for a specified destination",
    "parameters": {
        "type": "dict",
        "properties": {
            "origin": {
                "type": "string",
                "description": "The departure airport or city"
            "destination": {
                "type": "string",
                "description": "The destination airport or city"
            "departure_date": {
                "type": "string",
                "description": "The date of departure (in the format of YYYY-MM-DD)"
        "required": [
            "origin",
            "destination",
            "departure_date"
The current time is 2024-10-02 18:18:11. [User] I need to book a flight from New York to London for a business trip tomorrow. [Assistant] [[[[bookFlight(origin="New York", destination="London", departure_date="2024-10-03")]]]] [User] Please refine your answer. Directly output the refined answer, or the original answer if you think it is already perfect. [Assistant] [[[[bookFlight(origin="New York", destination="London", departure_date="2024-10-03")]]]]
Figure 7: One example for refine data. In this example, the assistant does not change the tool invocation as it is correct already. Through this kind of data, the model is able to learn when to stop iteration during inference.

### v2

@@
-ToolACE-R: Tool Learning with Adaptive Self-Refinement
+ToolACE-R: Model-aware Iterative Training and Adaptive Refinement for Tool Learning
 Abstract
 Tool learning, which allows Large Language Models (LLMs) to leverage external tools for solving complex user tasks, has emerged as a promising avenue for extending model capabilities.
-However, current approaches primarily focus on data synthesis for fine-tuning LLMs to invoke tools effectively, largely ignoring how to fully stimulate the potential of the model.
-In this paper, we propose ToolACE-R, a novel method that introduces adaptive self-refinement for tool invocations.
-Our approach features a model-aware iterative training procedure that progressively incorporates more training samples based on the model’s evolving capabilities.
-Additionally, it allows LLMs to iteratively refine their tool calls, optimizing performance without requiring external feedback.
-To further enhance computational efficiency, we integrate an adaptive mechanism when scaling the inference time, enabling the model to autonomously determine when to stop the refinement process.
-We conduct extensive experiments across several benchmark datasets, showing that ToolACE-R achieves competitive performance compared to advanced API-based models, even without any refinement.
-Furthermore, its performance can be further improved efficiently through adaptive self-refinement.
-Our results demonstrate the effectiveness of the proposed method, which is compatible with base models of various sizes, offering a promising direction for more efficient tool learning.
-ToolACE-R: Tool Learning with Adaptive Self-Refinement
-Xingshan Zeng1, Weiwen Liu1, Xu Huang2, Zezhong Wang3, Lingzhi Wang4, Liangyou Li1, Yasheng Wang1, Lifeng Shang1, Xin Jiang1, Ruiming Tang1, Qun Liu1 1Huawei Noah’s Ark Lab 2University of Science and Technology of China 3The Chinese University of Hong Kong 4Harbin Institute of Technology, Shenzhen zeng.xingshan@huawei.com, liuweiwen8@huawei.com, xuhuangcs@mail.ustc.edu.cn
+However, existing approaches primarily focus on data synthesis for fine-tuning LLMs to invoke tools effectively, largely ignoring how to fully stimulate the potential of the model.
+In this paper, we propose ToolACE-R, a novel framework that includes both model-aware iterative training and adaptive refinement for tool learning.
+ToolACE-R features a model-aware iterative training procedure that progressively adjust training samples based on the model’s evolving capabilities to maximize its potential.
+Additionally, it incorporates self-refinement training corpus which emphasizes LLM’s ability to iteratively refine their tool calls, optimizing performance without requiring external feedback.
+Furthermore, we introduce adaptive self-refinement mechanism for efficient test-time scaling, where the trained model can autonomously determine when to stop the process based on iterative self-refinement.
+We conduct extensive experiments across several benchmark datasets, showing that ToolACE-R achieves competitive performance compared to advanced API-based models.
+The performance of tool invocation can be further improved efficiently through adaptive self-refinement.
+These results highlight the effectiveness and generalizability of ToolACE-R, offering a promising direction for more efficient and scalable tool learning.
 1 Introduction
 Tool learning, which enables Large Language Models (LLMs) to leverage external tools to address complex user requirements, has gained increasing attention.
-With tool integration, LLMs can access up-to-date information, perform intricate computations, and utilize third-party services, significantly expanding their capabilities beyond simple natural language communication with humans Qu et al.
-(2024b).
-While tool invocation requires the LLMs to demonstrate strong understanding, reasoning, and instruction-following skills, customized fine-tuning is currently the dominant approach for enabling models to call external tools Qin et al.
-(2023); Patil et al.
-(2023); Liu et al.
-(2024b, c).
-Due to the limited availability of high-quality data, existing research has primarily focused on developing effective and efficient methods for data synthesis using advanced models Liu et al.
-(2024b, c); Wang et al.
-(2024b); Abdelaziz et al.
-(2024).
+With tool integration, LLMs can access up-to-date information, perform intricate computations, and utilize third-party services, significantly expanding their capabilities beyond simple natural language communication with humans [27].
+While tool invocation requires the LLMs to demonstrate strong understanding, reasoning, and instruction-following skills, customized fine-tuning is currently the dominant approach for enabling models to call external tools [19, 20, 23, 25].
+Due to the limited availability of high-quality data, existing research has primarily focused on developing effective and efficient methods for data synthesis using advanced models [1, 19, 20, 35].
 However, data synthesized by advanced models from different sources can lead to compatibility issues.
-Specifically, when synthesized samples exceed the model’s current knowledge, they may undermine the model’s performance or lead to hallucinations Ren et al.
-(2024); Kang et al.
-(2024).
+Specifically, when synthesized samples exceed the model’s current knowledge, they may undermine the model’s performance or lead to hallucinations [12, 29].
 As a result, determining which training data samples are appropriate for a given model remains an unresolved challenge.
-Additionally, the potential of maximizing the model’s own capabilities remains under exploration in the field of tool learning.
-It has been shown that a model’s reasoning performance can be improved by scaling test-time computation Brown et al.
-(2024); Snell et al.
-(2024).
-One promising approach involves iteratively refining the model’s output multiple times and then selecting the majority answer Snell et al.
-(2024).
-However, user queries vary in complexity, ranging from simple to highly complex.
-Scaling test-time computation for overly simple and straightforward questions is inefficient, as models can often answer them correctly without additional refinement.
-Therefore, an adaptive strategy for scaling test-time computation is needed to dynamically determine the appropriate level of effort required for each question.
-To address the challenges outlined above, we propose ToolACE-R, a method that enables adaptive self-refinement of tool invocations through a customized iterative training procedure.
-Specifically, we introduce a model-aware iterative training approach that facilitates LLMs in learning self-refinement for tool invocation in an iterative fashion, enhanced by model-based data selection.
-Additionally, by constructing refining data samples that preserve identical cases, LLMs are more likely to learn appropriate stopping criteria during the iterative process.
-Building on this training procedure, we further integrate adaptive self-refinement into iterative inference for inference time scaling, allowing models to autonomously determine when to halt the iteration, thereby improving scaling efficiency.
-We have conducted extensive experiments on several representative tool-calling benchmarks, such as the Berkeley Function Call Leaderboard (BFCL) Yan et al.
-(2024) and API-Bank Li et al.
-(2023), to assess the effectiveness and efficiency of ToolACE-R.The experimental results show that ToolACE-R achieves competitive performance compared to advanced API-based models like GPT-4o, and can be further enhanced efficiently through adaptive self-refinement.
-Additional analyses highlight the impact of the proposed modules and examine the generalizability of the method across different model backbones and sizes.
+Additionally, the potential for maximizing a model’s intrinsic capabilities remains underexplored in the field of tool learning.
+On one hand, prior work has seldom investigated how to fully exploit existing training data through techniques such as data augmentation.
+On the other hand, scaling test-time computation – despite its demonstrated effectiveness in enhancing LLM reasoning [4, 32] – has received limited attention in tool learning scenarios.
+A promising approach involves iteratively refining the model’s outputs and selecting the majority answer [32].
+However, user queries vary widely in complexity, ranging from simple to highly intricate.
+Scaling test-time computation indiscriminately is inefficient, as models can often correctly answer straightforward questions without the need for additional refinement.
+Therefore, an adaptive strategy for scaling test-time computation is needed to dynamically determine the appropriate level of computational effort required for each query.
+To address the challenges outlined above, we propose ToolACE-R, a novel pipeline for stimulating model potential in tool learning, encompassing both training and inference procedures.
+Specifically, we propose a model-aware iterative training framework that enables LLMs to learn tool invocation in alignment with their evolving capabilities.
+This is achieved through a novel model-aware difficulty metric, guiding the training process in an iterative manner.
+Additionally, a simple yet effective self-refinement data incorporation strategy is employed as a form of data augmentation.
+Furthermore, by preserving refinement samples that include identical cases, LLMs are encouraged to learn appropriate stopping criteria during the iterative process.
+Building upon this training procedure, we integrate adaptive self-refinement into iterative inference, allowing models to autonomously determine when to halt refinement, thereby improving inference-time efficiency.
+We have conducted extensive experiments on several representative tool-calling benchmarks, such as the Berkeley Function Call Leaderboard (BFCL) [38] and API-Bank [15], to assess the effectiveness and efficiency of ToolACE-R.
+Experimental results demonstrate that ToolACE-R achieves competitive performance compared to advanced API-based models such as GPT-4o, and can be further enhanced through adaptive self-refinement.
+Additional analyses highlight the contributions of the proposed modules and examine the generalizability of our method across different model backbones and sizes.
 The key contributions of this work are summarized as follows:
-  • We propose a model-aware iterative training procedure, enabling LLMs learning adaptive self-refinement for tool learning, which maximizes the model potential without any feedback.
-  • We integrate adaptive self-refinement into iterative inference, enabling models to autonomously decide when to halt the refinement process and allowing for more efficient use of computational resources when scaling test-time compute.
+  • We propose a model-aware iterative training procedure, augmented with self-refinement corpus construction, to maximize the potential of LLMs for tool learning.
+  • Building upon the training process, we integrate adaptive self-refinement into iterative inference, enabling models to autonomously decide when to halt the refinement process and allowing for more efficient use of computational resources when scaling test-time compute.
   • We conduct extensive experiments on several representative tool calling benchmarks, revealing effectiveness and efficiency of our method.
@@
 Tool Learning.
-Integrating external tools can enhance LLMs’ capabilities, enabling more specialized, precise, and reliable problem-solving for users (Qin et al., 2023).
 Tool learning methods generally fall into two categories, tuning-free and tuning-based approaches.
-Tuning-free methods employ prompting techniques, such as in-context tool-use examples, without additional training Mialon et al.
-(2023); Hsieh et al.
-(2023); Ruan et al.
-(2023).
-A notable example method is ReAct Yao et al.
-(2023), which can alternate between reasoning and action for complex tasks.
-Recent work Qu et al.
-(2024a) improves tool understanding by iteratively refining documentation.
-Tuning-based methods, which directly enhance tool calling ability via specialized fine-tuning, have recently gained much more attention Qin et al.
-(2023); Schick et al.
-(2023); Patil et al.
-(2023); Tang et al.
-(2023); Abdelaziz et al.
-(2024).
-Due to data scarcity, a large amount of them focus on data synthesis for improved generation procedures Liu et al.
-(2024b, c); Wang et al.
-(2024b), often overlooking the model’s inherent potential.
-While some others integrate tool feedback for refinement Du et al.
-(2024); Wang et al.
-(2024a), our approach is more streamlined, directly refining outputs in an iterative manner without external or any textual feedback to maximize model efficiency.
+Tuning-free methods employ prompting techniques without additional training [22, 8, 30].
+A notable example method is ReAct [40], which can alternate between reasoning and action for complex tasks.
+Recent work [26] improves tool understanding by iteratively refining documentation.
+Tuning-based methods, which directly enhance tool calling ability via specialized fine-tuning, have recently gained much more attention [25, 31, 23, 33, 1].
+A large amount of them focus on data synthesis for improved generation procedures [19, 20, 35], often overlooking the model’s inherent potential.
+While some others integrate tool feedback for refinement [7, 34], our approach is more streamlined, directly refining outputs in an iterative manner without external or any textual feedback to maximize model efficiency.
 Data Selection.
-Research indicates that selecting high-quality training samples is essential for fine-tuning LLMs Albalak et al.
-(2024).
-When the base model is sufficiently powerful, a small set of high-quality data can effectively harness the model’s potential, rather than relying on large quantities of general data Zhou et al.
-(2023); Liu et al.
-(2023).
-While earlier works primarily emphasize general data quality aspects, such as diversity and complexity, recent studies advocate for model-specific data selection Du et al.
-(2023); Li et al.
-(2024).
-This kind of approach is based on the observation that data distributions that significantly deviate from the base model’s are challenging for the model to learn from and may even degrade the performance Ren et al.
-(2024).
-Our method aligns with this perspective, proposing the use of pass@k metric to identify and select suitable training samples based on the model to be trained.
+Research indicates that selecting high-quality training samples is essential for fine-tuning LLMs [3].
+A small set of high-quality data can effectively harness the model’s potential, rather than relying on large quantities of data [42, 18].
+While earlier works primarily emphasize general data quality aspects, such as diversity and complexity, recent studies advocate for model-specific data selection [6, 14].
+This kind of approach is based on the observation that data distributions that significantly deviate from the base model’s are challenging for the model to learn from and may even degrade the performance [29].
+Our method aligns with this perspective, defining a metric with passrate@k to identify and select suitable training samples based on the model to be trained.
 Self Refinement.
-Previous research has demonstrated that LLMs can refine their own outputs through either self-feedback Madaan et al.
-(2023); Weng et al.
-(2023) or external feedback Qu et al.
-(2024c); Xu et al.
-(2024).
-Nevertheless, it remains challenging for LLMs to assess the correctness of their refined output autonomously Huang et al.
-(2024), which means that scaling test-time compute relying on iterative refinement still often depend on post-processing techniques, such as self-consistency Snell et al.
-(2024).
+Previous research has demonstrated that LLMs can refine their own outputs through either self-feedback [21, 36] or external feedback [28, 37].
+Nevertheless, it remains challenging for LLMs to assess the correctness of their refined output autonomously [10], which means that scaling test-time compute with iterative refinement still often depend on post-processing techniques, such as self-consistency [32].
 In this work, we propose a simple yet effective method that enables models to adaptively self-refine their outputs, thereby improving scaling efficiency.
-Figure 1: The overview of ToolACE-R, which includes two components, for training and inference, respectively.
+Figure 1: The overview of ToolACE-R, including two components, for training and inference, respectively.
 3 Methodology
@@
 The function $f_{\theta}(\cdot)$ denotes the generation process of the LLM with parameters $\theta$.
-The training sample for tool learning is typically formalized as a triplet consisting of a user query, candidate tools, and the ground-truth tool invocation: $\{\langle q,T\rangle,A\}$, where $q$ and $T$ serve as input and $A$ should be output.
+For each $a_{j}$, it might include several parameters and the corresponding values, denoted as $a_{j}=[p_{1}:v_{1},...,p_{i}:v_{i},$...], where $p_{i}$ is the parameter name, and $v_{i}$ is the corresponding value.
+The training samples for tool learning typically leverage $q$ and $T$ as context, and $A$ as output (the ground-truth tool invocation), denoted as $\{\langle q,T\rangle,A\}$.
 Our ToolACE-R is trained with a model-aware iterative training pipeline and can iteratively self-refine during inference in an adaptive manner. Figure 1 shows the overall procedure.
 3.2 Model-Aware Iterative Training
-Our model training starts with an instructed base model, ensuring that it has basic instruction following capability for our model-aware module.
-For training, we first gather a set of off-the-shelf training samples.
-Instead of using them directly, we apply a selection criterion to curate the data, which is then combined with generated self-refinement data.
-This curated and refined data is integrated into an iterative training process.
-Data Selection Criterion.
-Inspired by self-improvement techniques in existing LLM reasoning research Zelikman et al.
-(2022), where LLMs are trained using their own outputs, we introduce a pass@k criterion to select the most suitable training samples for each base model – those the model can answer (almost) correctly within at most $k$ tries.
-This approach may seem counterintuitive, as one might expect models to focus on learning from samples they struggle with.
-However, recent research Ren et al.
-(2024) has highlighted that models are more likely to perform self-alignment rather than learning new information during the fine-tuning stage.
-Based on this insight, we adopt this pass@k metric for data selection.
-Specifically, for each sample $\{\langle q,T\rangle,A\}$, the base model to be trained first generate the invocation up to $k$ times.
-If any of the responses approaches the groundtruth $A$, we retain it as an official training sample:
-$R(\langle q,T\rangle,A)=\begin{cases}1&\text{if }\exists\,j\in\{1,2,..,k\}\,\text{s.t. }\\ &f_{\theta_{0}}(\langle q,T\rangle)_{j}\approx A,\\ 0&\text{otherwise}.\end{cases}$ (2)
-Here $R(\langle q,T\rangle,A)=1$ indicates the training sample will be retained.
-$\theta_{0}$ represents the base model before training and $f_{\theta_{0}}(\cdot)_{j}$ means the $j$-th generation with the base model.
-A key aspect is how to define when a response "approaches" $A$.
-In our preliminary experiments, we found that if we restrict them to be exactly equal, the resulting training set still has an effect but limits the potential for performance improvement (see Table 4).
-However, by loosening this constraint, more suitable training samples are retained and the model’s potential can be more effectively harnessed.
-Therefore, we use the concept of "approaching" in this context.
-For any two invocations $A$ and $A^{\prime}$, we define $A\approx A^{\prime}$ if the used tools are exactly the same and each invocation only contains at most 1 different parameter values.
-Self-Refine Data Construction.
-To establish the model’s self-refinement capability, we additionally construct self-refine data based on the training samples.
-Specifically, we create a self-refine sample as a multi-turn interaction, where the first turn is the initial response from the model, and the second turn includes a refinement prompt from the user followed by the model’s refined answer: $\{\langle q,T\rangle,A_{1},r,A_{2}\}$.
-We apply the same selection criterion as described above, retaining the self-refinement training sample only when $A_{2}\approx A$:
-$R(\langle q,T\rangle,A_{1},r,A)=\begin{cases}1&\text{if }\exists\,j\in\{1,2,..,k\}\,\text{s.t. }\\ &f_{\theta_{0}}(\langle q,T\rangle,A_{1},r)_{j}\approx A,\\ 0&\text{otherwise}.\end{cases}$ (3)
-Here, we replace $A_{2}$ with $A$ to create the refined training sample: $\{\langle q,T\rangle,A_{1},r,A\}$.
+We begin model training with an instruction-tuned base model to ensure it possesses fundamental instruction-following capabilities required by our model-aware module.
+For the training process, we first collect a set of off-the-shelf samples and augment them via self-refinement.
+Rather than using these samples directly, we apply a selection criterion grounded in our defined notion of model-aware difficulty to curate the data.
+This curated dataset is then used to train the model and is iteratively re-curated based on new trained model throughout the training process.
+Model-Aware Difficulty.
+To guide tool learning with finer granularity, we introduce a model-aware difficulty metric that quantifies the inherent challenge a specific training sample poses to a given model.
+Rather than relying solely on static heuristics or dataset-level statistics, our definition measures how well the model can approximate the desired tool invocation behavior before training, thereby capturing a performance-aware notion of difficulty.
+Let a training sample be denoted as $\{\langle q,T\rangle,A\}$, where $A=[t_{1}(a_{1}),\dots,t_{m}(a_{m})]$ is the reference tool invocation sequence.
+Given a base model $\theta_{0}$, we obtain the model’s prediction as $A^{\prime}=[t^{\prime}_{1}(a^{\prime}_{1}),\dots,t^{\prime}_{m^{\prime}}(a^{\prime}_{m^{\prime}})]=f_{\theta_{0}}(\langle q,T\rangle)$.
+To assess the similarity between the predicted and reference sequences, we adopt Jaccard-style overlap to define a parameter-level alignment score between individual tool calls $t_{i}(a_{i})$ and $t^{\prime}_{j}(a^{\prime}_{j})$ as follows:
+$S_{ij}=\begin{cases}\frac{|a_{i}\cap a^{\prime}_{j}|}{|a_{i}\cup a^{\prime}_{j}|}&\text{if }t_{i}=t^{\prime}_{j}\\ 0&\text{otherwise }\end{cases}$ (2)
+Here, $|a_{i}\cap a^{\prime}_{j}|$ denotes the number of matching arguments between $a_{i}$ and $a^{\prime}_{j}$, i.e., parameters with both identical names and values.
+$|a_{i}\cup a^{\prime}_{j}|$ represents the total count of distinct arguments across both calls.
+Parameters with the same name but differing values are counted separately to penalize incorrect assignments.
+We aggregate all pairwise scores into a similarity matrix $S\in\mathbb{R}^{m\times m^{\prime}}$.
+To compute the best structural alignment between $A$ and $A^{\prime}$, we frame the task as a bipartite matching problem over the index sets $\mathcal{I}={1,\dots,m}$ and $\mathcal{J}={1,\dots,m^{\prime}}$, and apply the Hungarian algorithm [13] to identify the optimal matching $\mathcal{M}$ that maximizes the total alignment score:
+$\mathcal{M}={\arg\max}_{\mathcal{M}\in\text{Matchings}(\mathcal{I},\mathcal{J})}\sum_{(i,j)\in\mathcal{M}}S_{ij}$ (3)
+where $\text{Matchings}(\mathcal{I},\mathcal{J})$ denotes the set of all one-to-one matchings between $\mathcal{I}$ and $\mathcal{J}$, i.e., subsets of $\mathcal{I}\times\mathcal{J}$ in which each element in $\mathcal{I}$ and $\mathcal{J}$ appears in at most one pair.
+Using this optimal alignment, we define the overlap score between $A$ and $A^{\prime}$ as:
+$O(A,A^{\prime})=\frac{\sum_{(i,j)\in\mathcal{M}}S_{ij}}{\max(m,m^{\prime})}$ (4)
+This score reflects the degree to which the model reproduces the reference tool invocation, normalized by the longer of the two sequences to account for missing or superfluous tool calls.
+To incorporate model uncertainty and account for variance in its outputs, we evaluate the model over $k$ independent attempts, resulting in outputs $A^{\prime(1)},\dots,A^{\prime(k)}$}.
+The model-aware difficulty of a sample is then defined as:
+$D(\langle q,T\rangle,A)=1-\frac{1}{k}\sum_{l=1}^{k}O(A,A^{\prime(l)})$ (5)
+This difficulty score $D\in[0,1]$ quantifies the average divergence between the model’s output and the reference across multiple attempts.
+A lower score indicates that the model consistently produces accurate or nearly accurate tool calls, while a higher score suggests that the model struggles with the sample, failing to replicate even partial correctness.
+For instance, $D=0$ implies perfect performance across all outputs, whereas $D=1$ reflects complete failure in reproducing any reference calls.
+This model-aware difficulty serves as a valuable signal for iterative training, enabling more effective model development in tool-use scenarios.
+Data Augmentation with Self-Refinement.
+To augment the training corpus and establish the model’s self-refinement capability, we additionally construct self-refinement data based on the original training samples.
+Specifically, we create a self-refinement sample as a multi-turn interaction, where the first turn is the initial response from the model, and the second turn includes a refinement prompt from the user followed by the refined answer (typically the ground-truth tool invocation $A$): $\{\langle q,T\rangle,A_{1},r,A\}$.
+Here, the model to be trained will first produce its output tool invocation $A_{1}$.
+We then concatenate a refinement requirement prompt $r$ (a generalized sentence like “Please refine your answer…”) and the final ground-truth answer $A$.
 Notably, there may be samples where $A_{1}=A$, meaning the refined answer is identical to the previous answer.
 We retain these samples to enable the model to learn that "no further changes are needed when the answer is perfect." This is a crucial aspect of establishing the model’s adaptive, iterative self-refinement, and we will provide further details in Section 3.3.
-Iterative Training.
-Given the selected training samples (including self-refine samples), we apply supervised fine-tuning to the base model for one epoch.
-This constitutes one iteration.
-In subsequent iterations, we treat the trained model as the new base model, re-construct the self-refine samples, and re-select those that are suitable for fine-tuning the updated model.
-The model is then further trained using the newly selected samples.
-Ideally, with each iteration, the model’s performance improves, leading to an increase in the number of selected samples.
-The iteration process continues until the number of selected samples no longer increases, indicating that the model has saturated on this set of training samples.
+Each self-refinement sample can also be estimated its difficulty using Eq.
+5, by replacing $\{\langle q,T\rangle\}$ with $\{\langle q,T\rangle,A_{1},r\}$ as context.
+In this way, all the training samples including the constructed self-refinement samples can be estimated and labeled with a difficulty score.
+Model-Aware Iterative Data Selection and Training.
+We hypothesize that the most suitable training samples are those within the model’s current capability, either too difficult or too simple samples are ineffective for model training.
+To that end, we empirically set two thresholds $\alpha$ and $\beta$, where $\alpha<\beta$.
+For any training sample $\{C,A\}$ (where $C$ indicates the context, $\langle q,T\rangle$ or $\{\langle q,T\rangle,A_{1},r\}$ for self-refinement data), if its estimated difficulty score $\alpha<D(C,A)<\beta$, we retain it as an effective training sample and add to the actual training set $\mathcal{D}_{0}$.
+Given the selected training set $\mathcal{D}_{0}$, we apply supervised fine-tuning to the base model $\theta_{0}$ for one epoch, producing $\theta_{1}$.
+This constitutes the first iteration.
+In subsequent iterations, we treat the trained model as the new base model, re-construct the self-refinement samples, re-evaluate the difficulty, and re-select samples with the same criterion for fine-tuning the updated model.
+Ideally, with each iteration, the model’s performance improves, leading to an update of data difficulty and so new training samples might be included matching model’s current capability.
+The iteration will continue until the performance improvement is marginal, indicating that the model has saturated on this set of training samples.
 The resulting model will then be our final ToolACE-R model.
@@
 We propose an adaptive self-refine procedure that enables ToolACE-R to iteratively refine its own output and autonomously determine when to halt the iteration.
-The details are introduced in Algorithm 1.
-Specifically, given a user query $q$ and candidate tools $T$, ToolACE-R first generates the tool invocation $A_{0}$.
+Details are shown in Figure 1 (right hand side) and Algorithm 1 in Appendix.
+Specifically, given a user query $q$ and candidate tools $T$, ToolACE-R first generates the tool invocation $A^{\prime}$.
 In each subsequent iteration, ToolACE-R leverages the answer from the previous iteration as context and refines it.
@@
 To prevent infinite iteration, we set a maximum iteration limit $n$.
-Algorithm 1 Adaptive Self-Refine
-0:  Model $\theta$, query $q$, tools $T$, refine prompt $r$, time $n$
-0:  Refined final answer $A$
-1:  $A_{0}=f_{\theta}(\langle q,T\rangle)$  // First get the direct answer $A_{0}$
-2:  for $i=1$ to $n$ do
-3:     $A_{i}=f_{\theta}(\langle q,T\rangle,A_{i-1},r)$  // Get the refined answer
-4:     if $A_{i}==A_{i-1}$ then
-5:        return  $A_{i}$  // Adative stopping
-6:     end if
-7:  end for
-8:  return  $A_{n}$  // $n$ is the maximum iteration time
 Through adaptive self-refine, ToolACE-R is able to improve its performance, particularly when facing difficult cases.
 4 Experiments
+4.1 Experimental Settings
+Datasets and Models.
+We leverage a subset of ToolACE [19] training data as our original training data.
+The dataset contains multiple types of tool calling data, including cases with multiple candidate tools and parallel function callings, in both Chinese and English.
+We select LLaMA3.1-8B-Instruct [2] for base model in our main experiments.
+We also conduct experiments on other base models, including Qwen2.5-Instruct-series [39] (0.5B, 1.5B, 3B and 7B) and Mistral-7B-Instruct-v0.2 [11], to validate the generalizability of ToolACE-R.
+We compare with the state-of-the-art API-based models, including GPT-series [^1] https://chatgpt.com and Gemini-2.5-Pro [^2] https://gemini.google.com/, as well as open-source models like DeepSeek-V3 [17], Llama3.1-8B/70B-Instruct [2] and Qwen2.5-7B-Instruct [39].
+We also compare with fine-tuned tool calling models like ToolACE-8B [19], Hammer2.1-7B [16] and xLAM-2-8b-fc-r [41, 24].
+Benchmarks and Evaluation.
+We conduct experiments on several representative benchmarks, including the Berkeley Function Call Leaderboard (BFCL) [38], ACEBench [5], API-Bank [15] and ToolAlpaca [33].
+To minimize external interference and simplifying data preparation, we mainly evaluate performance on single-turn tool calling queries.
+For BFCL, we evaluate both the Non-live and Live subsets, which correspond to synthetic test cases and real-world scenarios, respectively.
+Each subset includes four categories: Simple, Multiple, Parallel, and Multiple Parallel.
+For ACEBench, we evaluate only the English normal category, excluding multi-turn cases.
+This includes the test samples in atom category and single-turn category.
+For API-Bank and ToolAlpaca, we focus on the first step tool calling samples, disregarding further tool or retrieval feedback.
+All performance is reported in terms of accuracy, where only cases with entirely correct tool calling are counted as correct.
+Implementation Details.
+Given resource constraints, we employ the parameter-efficient fine-tuning method LoRA [9].
+All model modules are configured for LoRA fine-tuning, with a rank of 16 and an alpha value of 32.
+Training is performed with a global batch size of 64 and a learning rate of $1\times 10^{-4}$, following a cosine learning rate schedule with a warmup ratio of 0.1.
+For difficulty estimation, we set the generation temperature to 1.0 to encourage diversity and allow the model to explore a broader range of outputs.
+Model generates 8 times (i.e.
+$k=8$) for each sample to ensure robustness.
+During data selection, we set $\alpha=0$ and $\beta=0.9$, which means that we discard samples that are extremely easy (difficulty score equals to 0, i.e., the model always answer fully correctly) or too difficult (most attempts are total incorrect).
+For the constructed refinement data, we apply a loss mask to the first assistant turn of the conversation, as it may involve an incorrect tool invocation, making it unsuitable for the model to learn from this turn.
+During evaluation, we adopt greedy search to ensure stability.
+Unless otherwise specified, , the maximum iteration time for adaptive self-refine is set to $n=5$ .
+4.2 Main Results
 Non-Live Live Overall
 Models Simple Multiple Parallel
+Multiple
 Parallel
+Simple Multiple Parallel
 Multiple
-Simple Multiple Parallel
 Parallel
-Multiple
 Non-live Live Overall
-GPT-4o-2024-11-20 79.42 95.50 94.00 83.50 84.88 79.77 87.50 75.00 88.10 80.75 84.43
-Gemini-1.5-Pro-002 78.33 93.50 92.5 90.00 82.95 77.40 87.50 83.33 88.58 78.68 83.63
+GPT-4o-2024-11-20 77.17 95.00 93.50 85.00 84.50 79.30 87.50 70.83 87.67 80.24 83.96
+GPT-4.1-2025-04-14 80.50 94.00 93.00 87.50 85.66 76.54 93.75 75.00 88.75 78.46 83.60
 GPT-4o-mini-2024-07-18 80.08 90.50 89.50 87.00 81.40 76.73 93.75 79.17 86.77 77.87 82.32
-GPT-3.5-Turbo-0125 77.92 93.50 67.00 53.00 80.62 78.63 75.00 58.33 72.85 78.61 75.73
-DeepSeek-V3 78.67 95.5 91.00 91.50 83.72 82.15 81.25 62.50 89.17 82.09 85.63
+Gemini-2.5-Pro-Exp-03-25 67.50 91.00 83.50 73.00 79.46 64.96 81.25 79.17 78.75 68.17 73.46
+DeepSeek-V3 78.67 95.50 91.00 91.50 83.72 82.15 81.25 62.50 89.17 82.09 85.63
 Llama3.1-70B-Inst 77.92 96.00 94.50 91.50 78.29 76.16 87.50 66.67 89.98 76.53 83.26
@@
 ToolACE-8B (FC) 76.67 93.50 90.50 89.50 73.26 76.73 81.25 70.83 87.54 76.02 81.78
-Hammer2.1-7B (FC) 78.08 95.00 93.50 88.00 76.74 77.4 81.25 70.83 88.65 77.20 82.92
-xLAM-7B-r (FC) 74.25 95.50 81.00 73.50 72.09 74.93 50.00 62.50 81.06 73.87 77.47
-ToolACE-R (FC) 79.92 96.50 93.50 93.00 81.78 81.01 87.50 83.33 90.73 81.27 86.00
-  + Adaptive Self-Refine 81.17 96.00 94.00 93.00 82.95 81.58 87.50 83.33 91.04 81.94 86.49
-Table 1: Accuracy comparison on BFCL (Last updated on 2025-01-18).
+Hammer2.1-7B (FC) 78.08 95.00 93.50 88.00 76.74 77.40 81.25 70.83 88.65 77.20 82.92
+xLAM-2-8b-fc-r (FC) 73.08 93.50 87.00 84.00 74.81 66.29 56.25 50.00 84.40 67.51 75.95
+ToolACE-R (FC) 79.75 97.00 95.50 91.50 82.95 81.10 93.75 87.50 90.94 81.72 86.33
+Table 1: Accuracy comparison on BFCL (Last updated on 2025-04-25).
 The best results in each category are marked in bold.
@@
 FC indicates the models are fine-tuned for function calling.
-4.1 Experimental Settings
-Datasets and Models.
-We leverage a subset of ToolACE Liu et al.
-(2024b) training data as our original full training dataset.
-The dataset contains multiple types of tool calling data, including cases with multiple candidate tools and parallel function callings, in both Chinese and English.
-For base model to be trained, we select LLaMA3.1-8B-Instruct AI@Meta (2024) for our main experiments.
-We also conduct experiments on other base models, including Qwen2.5-Instruct-series Yang et al.
-(2024) (0.5B, 1.5B, 3B and 7B) and Mistral-7B-Instruct-v0.2 Jiang et al.
-(2023), to validate the generalizability of ToolACE-R.
-We compare with the state-of-the-art API-based models, including GPT-series [^1] https://chatgpt.com and Gemini-1.5-Pro [^2] https://gemini.google.com/, as well as open-source models like DeepSeek-V3 Liu et al.
-(2024a), Llama3.1-8B/70B-Instruct AI@Meta (2024) and Qwen2.5-7B-Instruct Yang et al.
-(2024).
-We also compare with fine-tuned tool calling models like ToolACE-8B Liu et al.
-(2024b), Hammer2.1-7B Lin et al.
-(2024) and xLAM-7b-r Zhang et al.
-(2024).
-Benchmarks and Evaluation.
-We conduct experiments on several representative benchmarks, including the Berkeley Function Call Leaderboard (BFCL) Yan et al.
-(2024), ACEBench Chen et al.
-(2025), API-Bank Li et al.
-(2023) and ToolAlpaca Tang et al.
-(2023).
-To fully explore the model’s potential while minimizing external interference and simplifying data preparation, we mainly evaluate performance on single-turn tool calling queries.
-For BFCL, we evaluate both the Non-live and Live subsets, which correspond to synthetic test cases and real-world scenarios, respectively.
-Each subset includes four categories: Simple, Multiple, Parallel, and Parallel Multiple.
-Simple and Multiple cases both involve a single invoked tool, but in Multiple cases, there are multiple candidate tools.
-Parallel and Parallel Multiple cases require invoking multiple tools simultaneously.
-For ACEBench, we evaluate only the English normal category, excluding multi-turn cases.
-This includes the test samples in atom category and single-turn category.
-For API-Bank and ToolAlpaca, we focus on the first step tool calling samples, disregarding further tool or retrieval feedback.
-Performance is reported in terms of accuracy for all benchmarks.
-Implementation Details.
-Given resource constraints, we employ the parameter-efficient fine-tuning method LoRA Hu et al.
-(2022).
-All model modules are configured for LoRA fine-tuning, with a rank of 16 and an alpha value of 32.
-Training is performed with a global batch size of 64 and a learning rate of $1\times 10^{-4}$, following a cosine learning rate schedule with a warmup ratio of 0.1.
-For data selection, we set the generation temperature to 1.0 to encourage diversity and allow the model to explore a broader range of outputs.
-Each sample is generated up to 8 times (pass@8) to ensure potential.
-For the constructed refined data, we apply a loss mask to the first assistant turn of the conversation, as it may involve an incorrect tool invocation, making it unsuitable for the model to learn from this turn.
-During evaluation, we adopt greedy search to ensure stability and impose a maximum iteration limit of $n=5$.
-4.2 Main Results
 Results on BFCL.
@@
   • The performance gap between API-based and open-source models is minimal.
-Large open-source models like DeepSeek-V3 and Llama3.1-70B-Inst show comparable or even superior overall performance to GPT-4o and Gemini-1.5-Pro.
+Large open-source models like DeepSeek-V3 and Llama3.1-70B-Inst show comparable or even superior overall performance to GPT models and Gemini-2.5-Pro.
   • Specialized fine-tuned models significantly benefit from domain-specific fine-tuning, allowing smaller models (7/8B) to compete with larger general models like Llama3.1-70B-Inst.
-  • Our model, ToolACE-R which leverages our proposed iterative training procedure, achieves the best overall performance, surpassing both large API-based and open-source models.
-  • The adaptive self-refinement mechanism further enhances the performance of ToolACE-R, with larger improvement observed in more complex categories (Live), highlighting the effectiveness of our inference method.
+  • Our model, ToolACE-R, which leverages our proposed model-aware iterative training procedure, achieves the best overall performance, surpassing both large API-based and open-source models.
+Figure 2: Performance of different iterative self-refine methods.
+“Vanilla Self-Refine” always picks the final answer at each iteration.
+“Adaptive Self-Refine” is our proposed method, while “Self-Consistency” selects the majority answer during the iteration.
+To further assess the effectiveness of our adaptive self-refine method for iterative inference, we conducted experiments comparing with two other inference-time scaling approaches.
+The first, termed "Vanilla Self-Refine," always selects the last answer at each iteration as the final answer.
+The second method, "Self-Consistency at Itr n", chooses the majority answer from the answers collected during iterations 1 to n.
+The results of these inference methods on BFCL, are shown in Figure 2.
+As observed, the performance of "Vanilla Self-Refine" fluctuates across iterations, indicating that the model does not always refine the answer correctly.
+In contrast, our "Adaptive Self-Refine" method demonstrates more stability across iterations, suggesting that adjacent identical answers serve as a reliable signal for stopping further iterations.
+Our statistics also show that the average iteration time of our method on BFCL is $2.4$ (when maximum iteration time $n=5$), indicating that most cases terminate early thus achieving consistent performance.
+Our method achieves competitive performance compared to the strong baseline "Self-Consistency at Itr 10", further validating advantages.
 Models ACEBench APIBank ToolAlpaca
-GPT-4o-2024-11-20 87.00 77.16 83.87
-GPT-4o-mini-2024-07-18 81.63 73.10 83.87
+GPT-4o 87.00 77.16 83.87
+GPT-4o-mini 81.63 73.10 83.87
 Llama3.1-8B-Inst 50.63 60.41 79.03
 ToolACE-R 83.88 75.89 85.48
-  + Adaptive Sef-Refine 84.00 81.22 88.71
-Table 2: Accuracy comparison on ACEBench, API-Bank and ToolAlpaca benchmarks.
+  + Adaptive SR 84.00 81.22 88.71
+Table 2: Accuracy comparison on ACEBench, API-Bank and ToolAlpaca benchmarks. “SR” is short for “Self-Refine”.
+Models Non-live Live Overall
+ToolACE-R + Adaptive SR 91.32 82.34 86.83
+ToolACE-R 90.94 81.72 86.33
+  - Data Selection 88.96 81.13 85.04
+    - SR Data 88.75 80.09 84.42
+Llama3.1-8B-Inst (base) 84.37 69.28 76.83
+Table 3: Ablation Study of our proposed modules on BFCL. “SR” is short for “Self-Refine”. The modules are progressively removed.
 Results on More Benchmarks.
@@
 These results validate the potential of Adaptive Self-Refine to progressively refine tool invocation answers based solely on the model’s own capabilities, opening up new possibilities for scaling inference in tool learning scenarios.
+Ablation Study.
+To evaluate the contribution of each proposed module, we conduct an ablation study by progressively removing them.
+The corresponding results on the BFCL dataset are presented in Table 3.
+As shown, the additional inference cost introduced by adaptive self-refinement leads to performance gains in both categories.
+Further removing model-aware data selection and the constructed self-refinement data both result in a clear performance drop, with the absence of model-aware data selection causing the most significant degradation.
+These findings demonstrate the effectiveness of our proposed modules, especially for the model-aware part which mainly comes from the novel definition of data sample difficulty.
+4.3 Effects of Model-Aware Iterative Training
+In this subsection, we explore several different choices when applying model-aware iterative training, to validate the effectiveness of our current proposal.
+Choices of Selection Thresholds.
+In our main experiments, we set $\alpha=0$ and $\beta=0.9$ to discard samples that are either extremely simple or too difficult for model training.
+Here we experiment with more different choices of the two thresholds.
+We turn to another three kinds of choices, preserving simpler samples, preserving more difficult samples, or preserving samples in more medium difficulty.
+Table 5 displays the evaluation results on BFCL.
+As shown, setting $\alpha=-1$ and $\beta=0.9$, which includes even extremely simple cases (i.e., $\alpha=0$), leads to a slight performance drop while increasing training time due to the inclusion of a larger training corpus.
+In contrast, preserving only relatively difficult cases – by setting $\alpha=0.1$ and $\beta=2$ – results in a notable performance degradation across both categories, with an overall decrease of approximately 1 point.
+This suggests that samples beyond the model’s current capability may hinder effective learning and limit performance improvements based on its existing potential.
+We further evaluate a stricter setting ($\alpha=0.1,\beta=0.8$), which discards more relatively simple and difficult cases.
+This also results in performance drop, especially for Live category.
+The likely reason is that an excessive number of training samples are discarded (reducing the dataset to approximately half the size of our main setting) while retaining only a small portion of the harder examples.
+Thresholds Non-live Live Overall
+$\alpha=0,\beta=0.9$ (ToolACE-R) 90.94 81.72 86.33
+$\alpha=-1,\beta=0.9$ 90.73 81.27 86.00
+$\alpha=0.1,\beta=2$ 89.27 81.35 85.31
+$\alpha=0.1,\beta=0.8$ 90.83 80.38 85.61
+Table 4: Accuracy performance on BFCL when applying different threshods during model-aware data selection. We always select samples that their difficulty score satisfy $\alpha<D<\beta$.
 Models Non-live Live Overall
-Llama3.1-8B-Inst (Base) 84.37 69.28 76.83
-ToolACE-R (Full) 91.04 81.94 86.49
-  - Adaptive Self-Refine 90.73 81.27 86.00
-    - Iterative Training 88.12 80.24 84.18
-      - Data Selection 88.96 81.13 85.04
-        - Refine Data 88.75 80.09 84.42
-Table 3: Ablation Study of our proposed modules.
-Ablation Study.
-To investigate the impact of the proposed modules, we conduct an ablation study by progressively removing them.
-The corresponding evaluation results on the BFCL dataset are shown in Table 3.
-As observed, removing both adaptive self-refinement and iterative training leads to a noticeable performance degradation.
-However, further removing data selection does not result in a continued drop in performance.
-In fact, performance improves.
-This suggests that model-aware data selection should be combined with iterative training to fully leverage the entire training set, as the initial data selection (performed by the base model) typically only captures a small subset of the available data (mostly easy ones).
-Iterative selection is needed to capture more complex cases, as the trained model is improved.
-Finally, removing the constructed refine data also negatively impacts performance, validating that self-refine data acts as a form of data augmentation, also enhancing the model’s direct output.
-4.3 Effects of Iterative
-In this subsection, we examine the effects of our two iterative modules, i.e., model-aware iterative training and adaptive self-refine iterative inference.
-Figure 2: Performance of each iteration during training.
-“Amount” indicates the amount of training samples in that iteration with our data selection module.
-Iteration 0 means the model before training, i.e., Llama3.1-8B-Inst.
-Iterative Training.
-Figure 2 illustrates the results of the iterative training process, showing both the accuracy performance and the amount of training data selected at each iteration.
-Iteration 0 corresponds to the base model, Llama3.1-8B-Inst.
-As depicted, the quantity of selected samples increases monotonically across iterations, suggesting that the model’s performance on the training set improves, as the selected samples are those the model is more likely to answer correctly (as indicated by pass@k).
-On the other hand, accuracy improves rapidly during the first two iterations, reaching its peak at iteration 3.
-Subsequently, accuracy plateaus, with no further improvement in later iterations.
-This suggests that excessive training could lead to overfitting on the training set.
-Figure 3: Performance of different iterative self-refine methods.
-“Vanilla Self-Refine” always picks the final answer at each iteration.
-“Adaptive Self-Refine” is our proposed method, while “Self-Consistency” selects the majority answer during the iteration.
-Iterative Inference.
-To assess the effects of our adaptive self-refine method for iterative inference, we conducted additional experiments comparing it with two other inference-time scaling approaches.
-The first, termed "Vanilla Self-Refine," always selects the last answer at each iteration as the final answer.
-The second method, "Self-Consistency at Itr n", chooses the majority answer from the answers collected during iterations 1 to n.
-The results of these inference methods on BFCL, all based on the same model, are shown in Figure 3.
-As observed, the performance of "Vanilla Self-Refine" fluctuates across iterations, indicating that the model does not always refine the answer correctly.
-In contrast, our "Adaptive Self-Refine" method demonstrates more stability across iterations, suggesting that adjacent identical answers serve as a reliable signal for stopping further iterations.
-The "Self-Consistency at Itr n" method likely represents an upper bound for answer selection, with our method achieving competitive performance compared to this strong baseline.
-Furthermore, the performance of our method remains consistent across different iteration counts (3, 5, 10), likely because most cases terminate early, well before reaching the maximum iteration limit.
-This observation is supported by the average iteration time of 2.4, as indicated by our statistics on BFCL.
-4.4 Effects of Data Selection
-In this subsection, we investigate the effects of our model-aware data selection module.
-First, we explore the impact of the $k$ value when using pass@k for sample selection.
-Next, we examine how the use of data selected by different models affects model performance, highlighting the impact of model-aware selection.
-Models Direct Refine Amount
-Select with Pass@8
-  First Iteration 80.24 80.61 120K
-  Best Iteration 81.27 81.94 150K
-Select with Pass@1
-  First Iteration 78.90 78.53 94K
-  Best Iteration 79.57 80.01 143K
-Table 4: Accuracy performance on BFCL Live category when data selection with Pass@8 and Pass@1.
-“Direct” and “Refine” indicate the results with direct output and adaptive self-refine, respectively.
-“Amount” corresponds to the amount of selected training samples.
-Data Selection with Pass@k.
-In our data selection criterion, we retain training samples that the base model can answer (approximately) correctly within at most $k$ tries (pass@k).
-In this section, we examine the effects of varying $k$ values.
-Specifically, we report the evaluation results using pass@8 and pass@1 in Table 4.
-As shown, although the final number of selected samples is similar between the two methods (150K vs.
-143K), the accuracy performance for pass@8 is considerably higher.
-The performance gap is already pronounced in the first iteration, suggesting that a more lenient selection criterion plays a crucial role in maximizing model potential and shaping the model’s distribution.
-Notably, this effect does not diminish over additional iterations.
-Based on these findings, we select a larger $k$ value for our experiments.
-Model-aware Selection.
-In our data selection process, we propose selecting training samples in a model-aware manner, ensuring that a sample is chosen only when the base model can answer it (almost) correctly.
-This is based on the assumption that samples aligning with the model’s distribution are more suitable for training.
-To further assess the impact of such a model-aware approach, we conduct an additional experiment.
-Instead of using data selected by the base model Llama3.1-8B-Inst, we train Llama3.1-8B-Inst with data selected by Qwen2.5-3B-Inst and Qwen2.5-7B-Inst, as well as with the intersection of all three models’ selected data – i.e., only the samples that all three models perform well on are included in the dataset.
-The accuracy results of the trained models, along with the corresponding training data amounts, are presented in Table 5.
+$k=8$ (ToolACE-R) 90.94 81.72 86.33
+  + Adaptive SR 91.32 82.34 86.83
+$k=1$ 90.94 80.98 85.96
+  + Adaptive SR 90.71 82.01 86.36
+Table 5: Accuracy performance on BFCL when setting $k$ value in Eq. 5 as 88 and 11. “SR” is short for “Self-Refine”.
+Choices of $k$.
+In Eq.
+5, we define the difficulty score by allowing the model to make $k$ attempts, thereby better exploring its potential and producing a more robust estimate.
+To evaluate the impact of different $k$ values, we compare our main setting ($k=8$) with a reduced setting ($k=1$).
+The corresponding results are shown in Table 5.
+The model trained with $k=1$ exhibits a notable performance drop, particularly in the Live category, where difficult cases play a more critical role.
+Upon examining the final training set, we observe that approximately $10\%$ of the samples within the difficulty range $[0.8,0.9]$ under the $k=8$ setting are excluded when using $k=1$.
+This provides evidence that a larger $k$ value yields a more stable difficulty estimation and facilitates the selection of more appropriate training samples.
+Choices of Models.
+We further examine the “model-aware” part, where we use the base model to be trained to evaluate the difficulty of training samples.
+We conduct an additional experiment.
+Instead of using data selected by the base model Llama3.1-8B-Inst, we train Llama3.1-8B-Inst with data selected by Qwen2.5-3B-Inst and Qwen2.5-7B-Inst, as well as with the intersection of all three models’ selected data – i.e., only the samples that all three models perform similar (within the same difficulty scope) are included in the dataset.
+The results of the trained models on BFCL, along with the amounts of corresponding training data (in the first iteration), are presented in Table 6.
 As shown, performance drops when using data selected by other models, even when the data volume is larger.
@@
 This observation underscores the benefits of model-aware data selection.
-Models Accuracy Amount
-ToolACE-R 80.61 120K
-Data selected with Qwen2.5-3B 79.50 125K
-Data selected with Qwen2.5-7B 80.38 134K
-Intersection of Three Datasets 80.27 108K
-Table 5: Accuracy performance on BFCL Live category in the first interation when using different selected data to train Llama3.1-8B-Inst, along with the data amount.
-“Qwen2.5-3/7B” is short for Qwen2.5-3/7B-Inst.
+Models Live Overall Amount
+ToolACE-R 81.72 86.33 120K
+Select w/ Qwen2.5-3B 80.83 85.50 125K
+Select w/ Qwen2.5-7B 81.27 85.97 134K
+Intersection Data 81.37 86.08 108K
+Table 6: Accuracy performance on BFCL when using different selected data to train Llama3.1-8B-Inst, along with the data amount. “Qwen2.5-3/7B” is short for Qwen2.5-3/7B-Inst.
+Models Non-live Live Overall
+Qwen2.5-7B-Inst 84.38 72.17 78.28
+  + Original Data 88.29 78.39 83.34
+  + ToolACE-R 89.23 81.20 85.21
+Mistral-7B-Inst-v0.2 55.25 55.07 55.16
+  + Original Data 84.06 70.17 77.12
+  + ToolACE-R 87.98 80.01 84.00
+Table 7: Accuracy performance of using different backbones on BFCL.
+Iterative Training.
+To examine the training dynamics, we visualize the iterative training process by illustrating the trends in both accuracy and the amount of training data across iterations, as shown in Figure 4.
+Iteration 0 corresponds to the base model, LLaMA3.1-8B-Inst.
+As depicted, the number of selected training samples decreases monotonically with each iteration, primarily because more samples are assigned a difficulty score of 0 and consequently excluded from the training corpus.
+Meanwhile, accuracy improves rapidly during the first two iterations, peaking at iterations 2 and 3.
+Beyond this point, performance plateaus, with no further gains observed in subsequent iterations.
+This pattern suggests that excessive training may lead to overfitting on the filtered dataset.
+Therefore, to save the training cost, we recommend iterative training at most 3 times.
 4.4 Further analysis
 In this subsection, we examine the generalizability of our method by varying the base model and exploring the scaling effects across different model sizes.
-Models Non-live Live Overall
-Raw (Qwen2.5-7B-Inst) 84.38 72.17 78.28
-  + Direct Data 88.29 78.39 83.34
-  + ToolACE-R 89.23 81.20 85.21
-Raw (Mistral-7B-Inst-v0.2) 55.25 55.07 55.16
-  + Direct Data 84.06 70.17 77.12
-  + ToolACE-R 87.98 80.01 84.00
-Table 6: Accuracy performance of Different backbones on BFCL.
-“+ Direct Data” indicates the model directly trained with the full original set, while “+ ToolACE-R” is the result with our method, both iterative training and adaptive self-refine are applied.
+Figure 3: Accuracy performance and training amount of each iteration during training. Iteration 0 means the model before training.
+Figure 4: Accuracy Performance of Qwen2.5-Series. “Base” refers to the Qwen2.5-Instruct models, while “+ ToolACE-R” uses our method.
 Different Backbones.
@@
 Model Size Scaling.
-Figure 4: Performance of models in Qwen2.5-Series.
-“Base” refers to the Qwen2.5-Instruct models, while “Direct” and “Adaptive Self-Refine” represent the results of models trained using our method and evaluated with direct output and Adaptive Self-Refine, respectively.
 We further examine whether our method can be applied to models of varying sizes and whether performance improves as model size increases.
@@
 The results, shown in Figure 4, display both the original performance and the performance after applying ToolACE-R.
-The direct output of ToolACE-R consistently improves performance across all sizes, particularly for smaller models like 0.5B and 1.5B, and the Adaptive Self-Refine inference technique further enhances these results.
+Applying ToolACE-R consistently improves performance across all sizes, particularly for smaller models like 0.5B and 1.5B.
 This demonstrates the effectiveness of ToolACE-R across a range of model sizes.
@@
 In this paper, we introduce ToolACE-R, a tool learning technique designed to enable large language models (LLMs) to effectively leverage external tools.
-ToolACE-R consists of two key components: model-aware iterative training and adaptive self-refine iterative inference.
-Model-aware iterative training maximizes the model’s potential by iteratively selecting the most appropriate samples for training.
+ToolACE-R consists of two components: model-aware iterative training and adaptive self-refine iterative inference.
+Model-aware iterative training maximizes the model’s potential by iteratively selecting the most appropriate samples for training, with a novel definition of sample difficulty.
 Adaptive self-refine further enhances performance by allowing the model to autonomously determine when to stop refining its own answers during iterative inference.
@@
 Additionally, our analyses highlight the impact of the proposed modules on model performance.
-Limitations
-We summarize the limitations in the following two points:
-First, due to resource constraints, we conduct experiments only on models with up to 8B parameters, using LORA.
-As shown in the scaling performance results in Figure 4, we have not yet observed any bottlenecks preventing larger models from benefiting from our ToolACE-R method.
-Second, our focus is on single-turn tool-calling scenarios, excluding the multi-turn and multi-step cases for simplicity.
-However, we believe that single-turn cases are a fundamental aspect of tool learning.
-The significant improvements observed across several representative benchmarks demonstrate the effectiveness of our method.
-We plan to explore ToolACE-R in more complex scenarios in future work.
 Appendix A Appendix
-A.1 Detailed Introduction to Benchmarks
+A.1 Algorithm for Adaptive Self-refine
+Please refer to Alg. 1 for detailed process of adaptive self-refine strategy.
+Algorithm 1 Adaptive Self-Refine
+0: Model $\theta$, query $q$, tool candidates $T$, refine prompt $r$, maximum iteration time $n$
+0: Refined final answer AA
+[…]
+2:for $i=1$ to $n$ do
+3:  $A_{i}=f_{\theta}(\langle q,T\rangle,A_{i-1},r)$  // Get the refined answer based on the answer of last iteration
+4:  if $A_{i}==A_{i-1}$ then
+5:   return  $A_{i}$  // Adative stopping
+6:  end if
+7:end for
+8:return  $A_{n}$  // Achieves the maximum iteration time
+A.2 Detailed Introduction to Benchmarks
 Below we introduce the used benchmarks with more details:
-  • BFCL Yan et al.
-(2024): BFCL has progressed over time and is currently at version 3, which includes three main categories: non-live, live, and multi-turn.
+  • BFCL [38]: BFCL has progressed over time and is currently at version 3, which includes three main categories: non-live, live, and multi-turn.
 This paper focuses on the non-live and live categories, which primarily consist of Python-style tool call data.
-These are further subdivided into four types: Simple, Multiple, Parallel, and Parallel Multiple.
+These are further subdivided into four types: Simple, Multiple, Parallel, and Multiple Parallel.
     • Single: A single evaluation is the simplest and most frequently encountered format, where the user provides a single function document, triggering exactly one tool call.
@@
     • Parallel: A parallel test sample involves triggering multiple tool calls simultaneously in response to a single user query. The model must determine how many calls are needed.
-    • Parallel Multiple: The parallel multiple cases combine the concepts of both parallel and multiple.
+    • Multiple Parallel: The parallel multiple cases combine the concepts of both parallel and multiple.
 Here, the model is provided with several function documents, and each corresponding tool call may be invoked zero or more times.
-  • ACEBench Chen et al.
-(2025): ACEBench is a comprehensive tool-use benchmark that offers more detailed granularity.
+  • ACEBench [5]: ACEBench is a comprehensive tool-use benchmark that offers more detailed granularity.
 It is categorized into three main types: Normal, Special, and Agent.
@@
 The Single-turn category includes both single and parallel cases.
-  • API-Bank Li et al.
-(2023): API-Bank is a dialogue-style tool call dataset, consisting of two settings: Call and Retrieve + Call.
+  • API-Bank [15]: API-Bank is a dialogue-style tool call dataset, consisting of two settings: Call and Retrieve + Call.
 In this dataset, the model is tasked with invoking predefined local Python tools based on the user’s requirements in the dialogue.
@@
 In this work, we focus on the Call setting in our experiments.
-  • ToolAlpaca Tang et al.
-(2023): It introduces a multi-agent simulator that uses GPT-4 to simulate the return values of tools.
+  • ToolAlpaca [33]: It introduces a multi-agent simulator that uses GPT-4 to simulate the return values of tools.
 The model can make modifications and re-call the tool based on the returned values (e.g., error messages).
@@
 In our experiments, we does not leverage the returned values for refinement for consistency.
-A.2 Detailed Results in ACEBench
+A.3 Detailed Results in ACEBench
 Table 8 displays the detailed results in each sub-category of ACEBench.
@@
 Table 8: Detailed accuracy comparison on ACEBench.
-A.3 Prompts
+A.4 Limitations
+We summarize the limitations of this work in the following two points:
+First, due to resource constraints, we conduct experiments only on models with up to 8B parameters, using LORA, without applying our ToolACE-R method to larger LLMs.
+However, as shown in the scaling performance results in Figure 4, we have not yet observed any bottlenecks preventing larger models from benefiting from our ToolACE-R method.
+Second, we focus exclusively on single-turn tool-calling scenarios, leaving multi-turn cases for future work to maintain simplicity and stability.
+We argue that single-turn interactions are a fundamental component of tool learning.
+The substantial improvements observed across several representative benchmarks highlight the effectiveness of our approach.
+We plan to extend ToolACE-R to more complex, multi-turn scenarios in future work.
+A.5 Broader Impacts
+The development of ToolACE-R represents a step forward in enabling more autonomous and capable large language models (LLMs) through efficient and scalable tool learning.
+By reducing reliance on externally curated feedback and training with LORA in small LLMs, ToolACE-R lowers the barriers to deploying high-performing LLMs in resource-constrained environments.
+However, the growing autonomy of LLMs in generating and refining tool-use behavior also raises potential concerns.
+As models become better at self-directed decision-making, ensuring safety, accountability, and alignment with human values becomes increasingly important.
+Misuse of such systems, e.g., in automating high-stakes tasks without sufficient oversight, could lead to unintended consequences.
+Moreover, as LLMs become more proficient at tool interaction, there is a risk of amplifying biases or inaccuracies embedded in external tools or data sources.
+While ToolACE-R focuses on technical robustness and efficiency, further research should explore mechanisms for value alignment, transparency, and fail-safe design in autonomous tool-using agents.
+A.6 Prompts
 We show the detailed system prompt and refine prompt in Figure 5 and Figure 6, respectively.
@@
 For the training data, we diversify the system prompts to ensure better generalizability.
-A.4 Data example
+A.7 Data example
 Figure 7 is an example of our refine data. The model will only learn the last turn, with previous turns as context.
@@
 You should only return the tool call in tools call sections.
-If you decide to invoke any of the function(s), you MUST put it in the format of: [[[[func_name1(params_name1=params_value1, params_name2=params_value2,…), func_name2(params)]]]] You should not include any other text in the response.
-Here is a list of functions in JSON format that you can invoke: {candidate tools} {other information}
+If you decide to invoke any of the function(s), you MUST put it in the format of:
+[[func_name1(params_name1=value1, params_name2=value2,…), func_name2(params)]]
+You should not include any other text in the response. Here is a list of functions in JSON format that you can invoke:
+{candidate tools}
+{other information}
 Figure 5: The system prompt for tool calling.
@@
 Figure 6: The refine prompt for tool calling.
-[System] You are an expert in composing functions.
+[System]
+You are an expert in composing functions.
 You are given a question and a set of possible functions.
@@
 You should only return the tool call in tools call sections.
-If you decide to invoke any of the function(s), you MUST put it in the format of: [[[[func_name1(params_name1=params_value1, params_name2=params_value2,…), func_name2(params)]]]] You should not include any other text in the response.
-Here is a list of functions in JSON format that you can invoke:
+If you decide to invoke any of the function(s), you MUST put it in the format of:
+[[func_name1(params_name1=value1, params_name2=value2,…), func_name2(params)]]
+You should not include any other text in the response. Here is a list of functions in JSON format that you can invoke:
     "name": "bookFlight",
@@
                 "type": "string",
-                "description": "The date of departure (in the format of YYYY-MM-DD)"
+                "description": "The date of departure (in YYYY-$M$-DD)"
         "required": [
@@
 The current time is 2024-10-02 18:18:11.
-[User] I need to book a flight from New York to London for a business trip tomorrow.
-[Assistant] [[[[bookFlight(origin="New York", destination="London", departure_date="2024-10-03")]]]] [User] Please refine your answer.
-Directly output the refined answer, or the original answer if you think it is already perfect.
-[Assistant] [[[[bookFlight(origin="New York", destination="London", departure_date="2024-10-03")]]]]
+[User]
+I need to book a flight from New York to London for a business trip tomorrow.
+[Assistant]
+[[bookFlight(origin="New York", destination="London", departure_date="2024-10-03")]]
+[User]
+Please refine your answer. Directly output the refined answer, or the original answer if you think it is already perfect.
+[Assistant]
+[[bookFlight(origin="New York", destination="London", departure_date="2024-10-03")]]
 Figure 7: One example for refine data.

### v3

@@
 Additionally, it incorporates self-refinement training corpus which emphasizes LLM’s ability to iteratively refine their tool calls, optimizing performance without requiring external feedback.
-Furthermore, we introduce adaptive self-refinement mechanism for efficient test-time scaling, where the trained model can autonomously determine when to stop the process based on iterative self-refinement.
-We conduct extensive experiments across several benchmark datasets, showing that ToolACE-R achieves competitive performance compared to advanced API-based models.
-The performance of tool invocation can be further improved efficiently through adaptive self-refinement.
+Furthermore, we introduce adaptive self-refinement for efficient test-time scaling, where the trained model can autonomously determine when to stop the process based on iterative self-refinement.
+We conduct extensive experiments across several benchmark datasets, showing that ToolACE-R achieves competitive performance compared to advanced LLMs.
+The performance can be further improved efficiently through adaptive self-refinement.
 These results highlight the effectiveness and generalizability of ToolACE-R, offering a promising direction for more efficient and scalable tool learning.
-1 Introduction
-Tool learning, which enables Large Language Models (LLMs) to leverage external tools to address complex user requirements, has gained increasing attention.
-With tool integration, LLMs can access up-to-date information, perform intricate computations, and utilize third-party services, significantly expanding their capabilities beyond simple natural language communication with humans [27].
-While tool invocation requires the LLMs to demonstrate strong understanding, reasoning, and instruction-following skills, customized fine-tuning is currently the dominant approach for enabling models to call external tools [19, 20, 23, 25].
-Due to the limited availability of high-quality data, existing research has primarily focused on developing effective and efficient methods for data synthesis using advanced models [1, 19, 20, 35].
+Introduction
+Tool learning, enabling Large Language Models (LLMs) to leverage external tools to address complex user requirements, has gained increasing attention.
+With tool integration, LLMs can access up-to-date information, perform intricate computations, and utilize third-party services, significantly expanding their capabilities beyond simple natural language communication with humans (Qu et al.
+2024b).
+While tool invocation requires the LLMs to demonstrate strong understanding, reasoning, and instruction-following skills, customized fine-tuning is currently the dominant approach for enabling models to call external tools (Liu et al.
+2024b, c; Patil et al.
+2023; Qin et al.
+2023).
+Due to the limited availability of high-quality data, existing research has primarily focused on developing effective and efficient methods for data synthesis using advanced models (Abdelaziz et al.
+2024; Liu et al.
+2024b, c; Wang et al.
+2024b; Prabhakar et al.
+2025).
 However, data synthesized by advanced models from different sources can lead to compatibility issues.
-Specifically, when synthesized samples exceed the model’s current knowledge, they may undermine the model’s performance or lead to hallucinations [12, 29].
+Specifically, when synthesized samples exceed the model’s current knowledge, they may undermine the model’s performance or lead to hallucinations (Kang et al.
+2024; Ren et al.
+2024).
 As a result, determining which training data samples are appropriate for a given model remains an unresolved challenge.
@@
 On one hand, prior work has seldom investigated how to fully exploit existing training data through techniques such as data augmentation.
-On the other hand, scaling test-time computation – despite its demonstrated effectiveness in enhancing LLM reasoning [4, 32] – has received limited attention in tool learning scenarios.
-A promising approach involves iteratively refining the model’s outputs and selecting the majority answer [32].
+On the other hand, scaling test-time computation – despite its demonstrated effectiveness in enhancing LLM reasoning (Brown et al.
+2024; Snell et al.
+2024) – has received limited attention in tool learning scenarios.
+A promising approach involves iteratively refining the model’s outputs and selecting the majority answer (Snell et al.
+2024).
 However, user queries vary widely in complexity, ranging from simple to highly intricate.
@@
 Building upon this training procedure, we integrate adaptive self-refinement into iterative inference, allowing models to autonomously determine when to halt refinement, thereby improving inference-time efficiency.
-We have conducted extensive experiments on several representative tool-calling benchmarks, such as the Berkeley Function Call Leaderboard (BFCL) [38] and API-Bank [15], to assess the effectiveness and efficiency of ToolACE-R.
+We have conducted extensive experiments on several representative tool-calling benchmarks, such as the Berkeley Function Call Leaderboard (BFCL) (Yan et al.
+2024) and API-Bank (Li et al.
+2023), to assess the effectiveness and efficiency of ToolACE-R.
 Experimental results demonstrate that ToolACE-R achieves competitive performance compared to advanced API-based models such as GPT-4o, and can be further enhanced through adaptive self-refinement.
 Additional analyses highlight the contributions of the proposed modules and examine the generalizability of our method across different model backbones and sizes.
-The key contributions of this work are summarized as follows:
+The contributions of this work are summarized as follows:
   • We propose a model-aware iterative training procedure, augmented with self-refinement corpus construction, to maximize the potential of LLMs for tool learning.
@@
   • We conduct extensive experiments on several representative tool calling benchmarks, revealing effectiveness and efficiency of our method.
-2 Related Work
+Related Work
 Tool Learning.
 Tool learning methods generally fall into two categories, tuning-free and tuning-based approaches.
-Tuning-free methods employ prompting techniques without additional training [22, 8, 30].
-A notable example method is ReAct [40], which can alternate between reasoning and action for complex tasks.
-Recent work [26] improves tool understanding by iteratively refining documentation.
-Tuning-based methods, which directly enhance tool calling ability via specialized fine-tuning, have recently gained much more attention [25, 31, 23, 33, 1].
-A large amount of them focus on data synthesis for improved generation procedures [19, 20, 35], often overlooking the model’s inherent potential.
-While some others integrate tool feedback for refinement [7, 34], our approach is more streamlined, directly refining outputs in an iterative manner without external or any textual feedback to maximize model efficiency.
+Tuning-free methods employ prompting techniques without additional training (Mialon et al.
+2023; Hsieh et al.
+2023; Ruan et al.
+2023).
+A notable example method is ReAct (Yao et al.
+2023), which can alternate between reasoning and action for complex tasks.
+Recent work (Qu et al.
+2024a) improves tool understanding by iteratively refining documentation.
+Tuning-based methods, which directly enhance tool calling ability via specialized fine-tuning, have recently gained much more attention (Qin et al.
+2023; Schick et al.
+2023; Patil et al.
+2023; Tang et al.
+2023; Abdelaziz et al.
+2024).
+A large amount of them focus on data synthesis for improved generation procedures (Liu et al.
+2024b, c; Wang et al.
+2024b), often overlooking the model’s inherent potential.
+Zeng et al.
+(2025) also apply iterative training but based on DPO (Rafailov et al.
+2023) which introduces more complexity.
+While some others integrate tool feedback for refinement (Du, Wei, and Zhang 2024; Wang et al.
+2024a), our approach is more streamlined, directly refining outputs in an iterative manner without external or any textual feedback to maximize model efficiency.
 Data Selection.
-Research indicates that selecting high-quality training samples is essential for fine-tuning LLMs [3].
-A small set of high-quality data can effectively harness the model’s potential, rather than relying on large quantities of data [42, 18].
-While earlier works primarily emphasize general data quality aspects, such as diversity and complexity, recent studies advocate for model-specific data selection [6, 14].
-This kind of approach is based on the observation that data distributions that significantly deviate from the base model’s are challenging for the model to learn from and may even degrade the performance [29].
-Our method aligns with this perspective, defining a metric with passrate@k to identify and select suitable training samples based on the model to be trained.
+Selecting high-quality training samples is essential for fine-tuning LLMs (Albalak et al.
+2024).
+A small set of high-quality data can effectively harness the model’s potential, rather than relying on large quantities of data (Zhou et al.
+2023; Liu et al.
+2023).
+While earlier works primarily emphasize general data quality aspects, such as diversity and complexity, recent studies advocate for model-specific data selection (Du, Zong, and Zhang 2023; Li et al.
+2024).
+This kind of approach is based on the observation that data distributions that significantly deviate from the base model’s are challenging for the model to learn from and may even degrade the performance (Ren et al.
+2024).
+Our method aligns with this perspective, defining a new metric to identify and select suitable training samples based on the model to be trained.
 Self Refinement.
-Previous research has demonstrated that LLMs can refine their own outputs through either self-feedback [21, 36] or external feedback [28, 37].
-Nevertheless, it remains challenging for LLMs to assess the correctness of their refined output autonomously [10], which means that scaling test-time compute with iterative refinement still often depend on post-processing techniques, such as self-consistency [32].
-In this work, we propose a simple yet effective method that enables models to adaptively self-refine their outputs, thereby improving scaling efficiency.
+Previous research has demonstrated that LLMs can refine their own generations through either self-feedback (Madaan et al.
+2023; Weng et al.
+2023) or external feedback (Qu et al.
+2024c; Xu et al.
+2024).
+Nevertheless, it remains challenging for LLMs to assess the correctness of their refined output autonomously (Huang et al.
+2024), where scaling test-time compute with iterative refinement still often depends on post-processing techniques, such as majority voting (Snell et al.
+2024).
+In this work, we propose a simple yet effective method that enables LLMs to adaptively self-refine their outputs, improving scaling efficiency.
 Figure 1: The overview of ToolACE-R, including two components, for training and inference, respectively.
-3 Methodology
-3.1 Problem Formulation
-Tool learning aims to address the following problem: given a user query $q$ and a set of candidate tools $T=\{t_{1},t_{2},\dots,t_{n}\}$, the objective is to generate the correct tool invocations.
+Methodology
+Problem Formulation
+Given a user query $q$ and a set of candidate tools $T=\{t_{1},t_{2},\dots,t_{n}\}$, the objective of tool learning is to generate the correct tool invocations.
 This includes selecting the most appropriate tools and extracting the suitable parameter values.
@@
 The training samples for tool learning typically leverage $q$ and $T$ as context, and $A$ as output (the ground-truth tool invocation), denoted as $\{\langle q,T\rangle,A\}$.
-Our ToolACE-R is trained with a model-aware iterative training pipeline and can iteratively self-refine during inference in an adaptive manner. Figure 1 shows the overall procedure.
-3.2 Model-Aware Iterative Training
+ToolACE-R is trained with a model-aware iterative training pipeline and can iteratively self-refine during inference in an adaptive manner. Figure 1 shows the overall procedure.
+Model-Aware Iterative Training
 We begin model training with an instruction-tuned base model to ensure it possesses fundamental instruction-following capabilities required by our model-aware module.
@@
 Model-Aware Difficulty.
-To guide tool learning with finer granularity, we introduce a model-aware difficulty metric that quantifies the inherent challenge a specific training sample poses to a given model.
+We introduce a model-aware difficulty metric that quantifies the inherent challenge a specific training sample poses to a given model.
 Rather than relying solely on static heuristics or dataset-level statistics, our definition measures how well the model can approximate the desired tool invocation behavior before training, thereby capturing a performance-aware notion of difficulty.
@@
 We aggregate all pairwise scores into a similarity matrix $S\in\mathbb{R}^{m\times m^{\prime}}$.
-To compute the best structural alignment between $A$ and $A^{\prime}$, we frame the task as a bipartite matching problem over the index sets $\mathcal{I}={1,\dots,m}$ and $\mathcal{J}={1,\dots,m^{\prime}}$, and apply the Hungarian algorithm [13] to identify the optimal matching $\mathcal{M}$ that maximizes the total alignment score:
+To compute the best structural alignment between $A$ and $A^{\prime}$, we frame the task as a bipartite matching problem over the index sets $\mathcal{I}={1,\dots,m}$ and $\mathcal{J}={1,\dots,m^{\prime}}$, and apply the Hungarian algorithm (Kuhn 1955) to identify the optimal matching $\mathcal{M}$ that maximizes the total alignment score:
 $\mathcal{M}={\arg\max}_{\mathcal{M}\in\text{Matchings}(\mathcal{I},\mathcal{J})}\sum_{(i,j)\in\mathcal{M}}S_{ij}$ (3)
@@
 $O(A,A^{\prime})=\frac{\sum_{(i,j)\in\mathcal{M}}S_{ij}}{\max(m,m^{\prime})}$ (4)
-This score reflects the degree to which the model reproduces the reference tool invocation, normalized by the longer of the two sequences to account for missing or superfluous tool calls.
+This score reflects the degree to which the model reproduces the reference tool call, normalized by the longer of the two sequences to account for missing or superfluous tool calls.
 To incorporate model uncertainty and account for variance in its outputs, we evaluate the model over $k$ independent attempts, resulting in outputs $A^{\prime(1)},\dots,A^{\prime(k)}$}.
@@
 Notably, there may be samples where $A_{1}=A$, meaning the refined answer is identical to the previous answer.
-We retain these samples to enable the model to learn that "no further changes are needed when the answer is perfect." This is a crucial aspect of establishing the model’s adaptive, iterative self-refinement, and we will provide further details in Section 3.3.
+We retain these samples to enable the model to learn that ”no further changes are needed when the answer is perfect.” This is a crucial aspect of establishing the model’s adaptive, iterative self-refinement, and we will provide further details in later section.
 Each self-refinement sample can also be estimated its difficulty using Eq.
@@
 The resulting model will then be our final ToolACE-R model.
-3.3 Adaptive Self-Refine Iterative Inference
+Adaptive Self-Refine Iterative Inference
 With model-aware iterative training, our ToolACE-R model can directly generate tool invocations based on a user query and candidate tools, and refine its answer when necessary.
 We propose an adaptive self-refine procedure that enables ToolACE-R to iteratively refine its own output and autonomously determine when to halt the iteration.
-Details are shown in Figure 1 (right hand side) and Algorithm 1 in Appendix.
+Details are shown in Figure 1 (right hand side).
 Specifically, given a user query $q$ and candidate tools $T$, ToolACE-R first generates the tool invocation $A^{\prime}$.
@@
 Therefore, the iteration continues until the answers from two consecutive iterations are identical.
-We refer to this procedure as adaptive self-refine, where the model itself determines when to stop the iterations.
+We refer to this procedure as adaptive self-refinement, where the model itself determines when to stop the iterations.
 To prevent infinite iteration, we set a maximum iteration limit $n$.
-Through adaptive self-refine, ToolACE-R is able to improve its performance, particularly when facing difficult cases.
-4 Experiments
-4.1 Experimental Settings
+Through adaptive self-refinement, ToolACE-R is able to improve its performance, particularly when facing difficult cases.
+Experiments
+Experimental Settings
 Datasets and Models.
-We leverage a subset of ToolACE [19] training data as our original training data.
-The dataset contains multiple types of tool calling data, including cases with multiple candidate tools and parallel function callings, in both Chinese and English.
-We select LLaMA3.1-8B-Instruct [2] for base model in our main experiments.
-We also conduct experiments on other base models, including Qwen2.5-Instruct-series [39] (0.5B, 1.5B, 3B and 7B) and Mistral-7B-Instruct-v0.2 [11], to validate the generalizability of ToolACE-R.
-We compare with the state-of-the-art API-based models, including GPT-series [^1] https://chatgpt.com and Gemini-2.5-Pro [^2] https://gemini.google.com/, as well as open-source models like DeepSeek-V3 [17], Llama3.1-8B/70B-Instruct [2] and Qwen2.5-7B-Instruct [39].
-We also compare with fine-tuned tool calling models like ToolACE-8B [19], Hammer2.1-7B [16] and xLAM-2-8b-fc-r [41, 24].
+We use a subset of ToolACE (Liu et al.
+2024b) training data as our training data, where we only retain the samples with single-turn tool calling for simplicity.
+We select LLaMA3.1-8B-Instruct (AI@Meta 2024) as base model in our main experiments.
+We also conduct experiments on Qwen2.5-Instruct-series (Yang et al.
+2024) (0.5B, 1.5B, 3B and 7B) and Mistral-7B-Instruct-v0.2 (Jiang et al.
+2023), to validate the generalizability of ToolACE-R.
+We compare with the state-of-the-art API-based models, including GPT-series and Gemini-2.5-Pro, as well as open-source models like DeepSeek-V3 (Liu et al.
+2024a), Llama3.1-8B/70B-Instruct (AI@Meta 2024) and Qwen2.5-7B-Instruct (Yang et al.
+2024).
+We also compare with fine-tuned tool calling models like ToolACE-8B (Liu et al.
+2024b), Hammer2.1-7B (Lin et al.
+2024) and xLAM-2-8b-fc-r (Zhang et al.
+2024; Prabhakar et al.
+2025).
 Benchmarks and Evaluation.
-We conduct experiments on several representative benchmarks, including the Berkeley Function Call Leaderboard (BFCL) [38], ACEBench [5], API-Bank [15] and ToolAlpaca [33].
+We conduct experiments on several representative benchmarks, including the Berkeley Function Call Leaderboard (BFCL) (Yan et al.
+2024), ACEBench (Chen et al.
+2025), API-Bank (Li et al.
+2023) and ToolAlpaca (Tang et al.
+2023).
 To minimize external interference and simplifying data preparation, we mainly evaluate performance on single-turn tool calling queries.
-For BFCL, we evaluate both the Non-live and Live subsets, which correspond to synthetic test cases and real-world scenarios, respectively.
+For BFCL, we evaluate the Non-live and Live subsets, which correspond to synthetic test cases and real-world scenarios, respectively.
 Each subset includes four categories: Simple, Multiple, Parallel, and Multiple Parallel.
@@
 All performance is reported in terms of accuracy, where only cases with entirely correct tool calling are counted as correct.
-Implementation Details.
-Given resource constraints, we employ the parameter-efficient fine-tuning method LoRA [9].
-All model modules are configured for LoRA fine-tuning, with a rank of 16 and an alpha value of 32.
-Training is performed with a global batch size of 64 and a learning rate of $1\times 10^{-4}$, following a cosine learning rate schedule with a warmup ratio of 0.1.
-For difficulty estimation, we set the generation temperature to 1.0 to encourage diversity and allow the model to explore a broader range of outputs.
-Model generates 8 times (i.e.
-$k=8$) for each sample to ensure robustness.
-During data selection, we set $\alpha=0$ and $\beta=0.9$, which means that we discard samples that are extremely easy (difficulty score equals to 0, i.e., the model always answer fully correctly) or too difficult (most attempts are total incorrect).
-For the constructed refinement data, we apply a loss mask to the first assistant turn of the conversation, as it may involve an incorrect tool invocation, making it unsuitable for the model to learn from this turn.
-During evaluation, we adopt greedy search to ensure stability.
-Unless otherwise specified, , the maximum iteration time for adaptive self-refine is set to $n=5$ .
-4.2 Main Results
 Non-Live Live Overall
@@
 FC indicates the models are fine-tuned for function calling.
+Implementation Details.
+We employ parameter-efficient fine-tuning method LoRA (Hu et al.
+2022) given resource constraints.
+All model modules are configured for LoRA fine-tuning, with a rank of 16 and an alpha value of 32.
+Training is performed with a global batch size of 64 and a learning rate of $1\times 10^{-4}$, following a cosine learning rate schedule with a warmup ratio of 0.1.
+For difficulty estimation, we set the generation temperature to 1.0 to encourage diversity and allow the model to explore a broader range of outputs.
+Model generates 8 times (i.e.
+$k=8$) for each sample to ensure robustness.
+During data selection, we set $\alpha=0$ and $\beta=0.9$, which means that we discard samples that are extremely easy (difficulty score equals to 0, i.e., the model always answer fully correctly) or too difficult (most attempts are total incorrect).
+For the constructed refinement data, we apply a loss mask to the first assistant turn of the conversation, as it may involve an incorrect tool invocation, making it unsuitable for the model to learn from this turn.
+During evaluation, we adopt greedy search to ensure stability.
+Unless otherwise specified, , the maximum iteration time for adaptive self-refinement is set to $n=5$ .
+Main Results
 Results on BFCL.
-Table 1 presents the comparison results on BFCL, including detailed results for each category and the overall average performance. The following key observations can be made:
+Table 1 presents the comparison results on BFCL, including detailed results for each category. The following key observations can be made:
   • The performance gap between API-based and open-source models is minimal.
-Large open-source models like DeepSeek-V3 and Llama3.1-70B-Inst show comparable or even superior overall performance to GPT models and Gemini-2.5-Pro.
-  • Specialized fine-tuned models significantly benefit from domain-specific fine-tuning, allowing smaller models (7/8B) to compete with larger general models like Llama3.1-70B-Inst.
+Large open-source models like DeepSeek-V3 show comparable or even superior overall performance to GPT models and Gemini-2.5-Pro.
+  • Specialized fine-tuned models significantly benefit from domain-specific fine-tuning, where smaller models (7/8B) are allowed to compete with larger general models like Llama3.1-70B-Inst.
   • Our model, ToolACE-R, which leverages our proposed model-aware iterative training procedure, achieves the best overall performance, surpassing both large API-based and open-source models.
@@
 To further assess the effectiveness of our adaptive self-refine method for iterative inference, we conducted experiments comparing with two other inference-time scaling approaches.
-The first, termed "Vanilla Self-Refine," always selects the last answer at each iteration as the final answer.
-The second method, "Self-Consistency at Itr n", chooses the majority answer from the answers collected during iterations 1 to n.
+The first, termed ”Vanilla Self-Refine,” always selects the last answer at each iteration as the final answer.
+The second method, ”Self-Consistency at Itr n”, chooses the majority answer from the answers collected during iterations 1 to n.
 The results of these inference methods on BFCL, are shown in Figure 2.
-As observed, the performance of "Vanilla Self-Refine" fluctuates across iterations, indicating that the model does not always refine the answer correctly.
-In contrast, our "Adaptive Self-Refine" method demonstrates more stability across iterations, suggesting that adjacent identical answers serve as a reliable signal for stopping further iterations.
+As observed, the performance of ”Vanilla Self-Refine” fluctuates across iterations, indicating that the model does not always refine the answer correctly.
+In contrast, our ”Adaptive Self-Refine” method demonstrates more stability across iterations, suggesting that adjacent identical answers serve as a reliable signal for stopping further iterations.
 Our statistics also show that the average iteration time of our method on BFCL is $2.4$ (when maximum iteration time $n=5$), indicating that most cases terminate early thus achieving consistent performance.
-Our method achieves competitive performance compared to the strong baseline "Self-Consistency at Itr 10", further validating advantages.
+Our method achieves competitive performance compared to the strong baseline ”Self-Consistency at Itr 10”, further validating advantages.
 Models ACEBench APIBank ToolAlpaca
@@
   + Adaptive SR 84.00 81.22 88.71
-Table 2: Accuracy comparison on ACEBench, API-Bank and ToolAlpaca benchmarks. “SR” is short for “Self-Refine”.
+Table 2: Accuracy comparison on ACEBench, API-Bank and ToolAlpaca. “SR” is short for “Self-Refine”.
 Models Non-live Live Overall
@@
 Llama3.1-8B-Inst (base) 84.37 69.28 76.83
-Table 3: Ablation Study of our proposed modules on BFCL. “SR” is short for “Self-Refine”. The modules are progressively removed.
+Table 3: Ablation Study of our proposed modules on BFCL. “SR” is short for “Self-Refine”.
 Results on More Benchmarks.
@@
 These findings demonstrate the effectiveness of our proposed modules, especially for the model-aware part which mainly comes from the novel definition of data sample difficulty.
-4.3 Effects of Model-Aware Iterative Training
+Effects of Model-Aware Iterative Training
 In this subsection, we explore several different choices when applying model-aware iterative training, to validate the effectiveness of our current proposal.
@@
 As shown, setting $\alpha=-1$ and $\beta=0.9$, which includes even extremely simple cases (i.e., $\alpha=0$), leads to a slight performance drop while increasing training time due to the inclusion of a larger training corpus.
-In contrast, preserving only relatively difficult cases – by setting $\alpha=0.1$ and $\beta=2$ – results in a notable performance degradation across both categories, with an overall decrease of approximately 1 point.
+In contrast, preserving only relatively difficult cases (by setting $\alpha=0.1$ and $\beta=2$) results in a notable performance degradation across both categories, with an overall decrease of approximately 1%.
 This suggests that samples beyond the model’s current capability may hinder effective learning and limit performance improvements based on its existing potential.
-We further evaluate a stricter setting ($\alpha=0.1,\beta=0.8$), which discards more relatively simple and difficult cases.
+We further evaluate a stricter setting ($\alpha=0.1,\beta=0.8$), which discards more simple and difficult cases.
 This also results in performance drop, especially for Live category.
@@
 $\alpha=0.1,\beta=0.8$ 90.83 80.38 85.61
-Table 4: Accuracy performance on BFCL when applying different threshods during model-aware data selection. We always select samples that their difficulty score satisfy $\alpha<D<\beta$.
+Table 4: Performance on BFCL when applying different threshods during model-aware data selection. We always select samples that their difficulty score satisfy $\alpha<D<\beta$.
 Models Non-live Live Overall
@@
   + Adaptive SR 90.71 82.01 86.36
-Table 5: Accuracy performance on BFCL when setting $k$ value in Eq. 5 as 88 and 11. “SR” is short for “Self-Refine”.
+Table 5: Performance on BFCL when setting $k$ value in Eq. 5 as 88 and 11. “SR” is short for “Self-Refine”.
 Choices of $k$.
@@
 Intersection Data 81.37 86.08 108K
-Table 6: Accuracy performance on BFCL when using different selected data to train Llama3.1-8B-Inst, along with the data amount. “Qwen2.5-3/7B” is short for Qwen2.5-3/7B-Inst.
+Table 6: Performance on BFCL when using different selected data to train Llama3.1-8B-Inst, along with the data amount. “Qwen2.5-3/7B” is short for Qwen2.5-3/7B-Inst.
 Models Non-live Live Overall
@@
   + ToolACE-R 87.98 80.01 84.00
-Table 7: Accuracy performance of using different backbones on BFCL.
+Table 7: Performance of using different models on BFCL.
 Iterative Training.
@@
 Therefore, to save the training cost, we recommend iterative training at most 3 times.
-4.4 Further analysis
+Further analysis
 In this subsection, we examine the generalizability of our method by varying the base model and exploring the scaling effects across different model sizes.
-Figure 3: Accuracy performance and training amount of each iteration during training. Iteration 0 means the model before training.
-Figure 4: Accuracy Performance of Qwen2.5-Series. “Base” refers to the Qwen2.5-Instruct models, while “+ ToolACE-R” uses our method.
+Figure 3: Performance and training amount of each iteration during training. Iteration 0 means the model before training.
+Figure 4: Performance of Qwen2.5-Series. “Base” refers to the Qwen2.5-Instuct models without any training, while “+ ToolACE-R” uses our method.
 Different Backbones.
@@
 We further examine whether our method can be applied to models of varying sizes and whether performance improves as model size increases.
-For this, we use Qwen2.5-Instruct models of different sizes, including 0.5B, 1.5B, 3B, and 7B, as base models.
+We use Qwen2.5-Instruct models of different sizes, including 0.5B, 1.5B, 3B, and 7B, as base models.
 The results, shown in Figure 4, display both the original performance and the performance after applying ToolACE-R.
@@
 This demonstrates the effectiveness of ToolACE-R across a range of model sizes.
-5 Conclusion
-In this paper, we introduce ToolACE-R, a tool learning technique designed to enable large language models (LLMs) to effectively leverage external tools.
+Conclusion
+In this paper, we introduce ToolACE-R, a tool learning technique designed to enable LLMs to learn to effectively leverage external tools.
 ToolACE-R consists of two components: model-aware iterative training and adaptive self-refine iterative inference.
 Model-aware iterative training maximizes the model’s potential by iteratively selecting the most appropriate samples for training, with a novel definition of sample difficulty.
-Adaptive self-refine further enhances performance by allowing the model to autonomously determine when to stop refining its own answers during iterative inference.
-Experimental results demonstrate that ToolACE-R consistently improves performance across various tool-calling benchmarks and is effective with different backbone models.
+Adaptive self-refinement further enhances performance by allowing the model to autonomously determine when to stop refining its own answers during iterative inference.
+Experiments demonstrate that ToolACE-R consistently improves performance across various benchmarks and is effective with different backbone models.
 Additionally, our analyses highlight the impact of the proposed modules on model performance.
-Appendix A Appendix
-A.1 Algorithm for Adaptive Self-refine
-Please refer to Alg. 1 for detailed process of adaptive self-refine strategy.
-Algorithm 1 Adaptive Self-Refine
-0: Model $\theta$, query $q$, tool candidates $T$, refine prompt $r$, maximum iteration time $n$
-0: Refined final answer AA
-[…]
-2:for $i=1$ to $n$ do
-3:  $A_{i}=f_{\theta}(\langle q,T\rangle,A_{i-1},r)$  // Get the refined answer based on the answer of last iteration
-4:  if $A_{i}==A_{i-1}$ then
-5:   return  $A_{i}$  // Adative stopping
-6:  end if
-7:end for
-8:return  $A_{n}$  // Achieves the maximum iteration time
-A.2 Detailed Introduction to Benchmarks
-Below we introduce the used benchmarks with more details:
-  • BFCL [38]: BFCL has progressed over time and is currently at version 3, which includes three main categories: non-live, live, and multi-turn.
-This paper focuses on the non-live and live categories, which primarily consist of Python-style tool call data.
-These are further subdivided into four types: Simple, Multiple, Parallel, and Multiple Parallel.
-    • Single: A single evaluation is the simplest and most frequently encountered format, where the user provides a single function document, triggering exactly one tool call.
-    • Multiple: The multiple category involves a user query that selects a single tool call from 2 to 4 available function documents.
-The model must choose the most appropriate tool to call based on the user’s context.
-    • Parallel: A parallel test sample involves triggering multiple tool calls simultaneously in response to a single user query. The model must determine how many calls are needed.
-    • Multiple Parallel: The parallel multiple cases combine the concepts of both parallel and multiple.
-Here, the model is provided with several function documents, and each corresponding tool call may be invoked zero or more times.
-  • ACEBench [5]: ACEBench is a comprehensive tool-use benchmark that offers more detailed granularity.
-It is categorized into three main types: Normal, Special, and Agent.
-In this work, we focus on the Atom and Single-turn sub-categories within the Normal category.
-Atom cases involve a set of APIs that contain specific parameter types, such as booleans, enumerations, numbers, lists, and objects.
-The Single-turn category includes both single and parallel cases.
-  • API-Bank [15]: API-Bank is a dialogue-style tool call dataset, consisting of two settings: Call and Retrieve + Call.
-In this dataset, the model is tasked with invoking predefined local Python tools based on the user’s requirements in the dialogue.
-Accuracy is assessed by comparing whether the tool’s returned values match the ground truth.
-In this work, we focus on the Call setting in our experiments.
-  • ToolAlpaca [33]: It introduces a multi-agent simulator that uses GPT-4 to simulate the return values of tools.
-The model can make modifications and re-call the tool based on the returned values (e.g., error messages).
-Ultimately, GPT-4 is used to evaluate the accuracy.
-In our experiments, we does not leverage the returned values for refinement for consistency.
-A.3 Detailed Results in ACEBench
-Table 8 displays the detailed results in each sub-category of ACEBench.
-As can be seen, in all sub-categories, ToolACE-R performs much better than its base model Llama3.1-8B-Inst.
-The improvement of Adaptive Self-Refine is not significant.
-Atom Single-Turn Overall
-Models Bool Enum Number List
-Object
-Short
-Object
-Deep
-Single Parallel Overall
-GPT-4o-2024-11-20 94.00 94.00 98.00 94.00 64.00 96.00 84.00 72.00 87.00
-GPT-4o-mini-2024-07-18 88.00 94.00 98.00 90.00 44.00 92.00 77.00 70.00 81.63
-Llama3.1-8B-Inst 50.00 74.00 48.00 60.00 30.00 44.00 55.00 44.00 50.63
-ToolACE-R 88.00 96.00 94.00 94.00 50.00 92.00 83.00 74.00 83.88
-  + Adaptive Self-Refine 88.00 96.00 94.00 94.00 50.00 94.00 84.00 72.00 84.00
-Table 8: Detailed accuracy comparison on ACEBench.
-A.4 Limitations
-We summarize the limitations of this work in the following two points:
-First, due to resource constraints, we conduct experiments only on models with up to 8B parameters, using LORA, without applying our ToolACE-R method to larger LLMs.
-However, as shown in the scaling performance results in Figure 4, we have not yet observed any bottlenecks preventing larger models from benefiting from our ToolACE-R method.
-Second, we focus exclusively on single-turn tool-calling scenarios, leaving multi-turn cases for future work to maintain simplicity and stability.
-We argue that single-turn interactions are a fundamental component of tool learning.
-The substantial improvements observed across several representative benchmarks highlight the effectiveness of our approach.
-We plan to extend ToolACE-R to more complex, multi-turn scenarios in future work.
-A.5 Broader Impacts
-The development of ToolACE-R represents a step forward in enabling more autonomous and capable large language models (LLMs) through efficient and scalable tool learning.
-By reducing reliance on externally curated feedback and training with LORA in small LLMs, ToolACE-R lowers the barriers to deploying high-performing LLMs in resource-constrained environments.
-However, the growing autonomy of LLMs in generating and refining tool-use behavior also raises potential concerns.
-As models become better at self-directed decision-making, ensuring safety, accountability, and alignment with human values becomes increasingly important.
-Misuse of such systems, e.g., in automating high-stakes tasks without sufficient oversight, could lead to unintended consequences.
-Moreover, as LLMs become more proficient at tool interaction, there is a risk of amplifying biases or inaccuracies embedded in external tools or data sources.
-While ToolACE-R focuses on technical robustness and efficiency, further research should explore mechanisms for value alignment, transparency, and fail-safe design in autonomous tool-using agents.
-A.6 Prompts
-We show the detailed system prompt and refine prompt in Figure 5 and Figure 6, respectively.
-Note that the displayed prompts are those used during inference.
-For the training data, we diversify the system prompts to ensure better generalizability.
-A.7 Data example
-Figure 7 is an example of our refine data. The model will only learn the last turn, with previous turns as context.
-You are an expert in composing functions.
-You are given a question and a set of possible functions.
-Based on the question, you will need to make one or more function/tool calls to achieve the purpose.
-If none of the function can be used, point it out.
-If the given question lacks the parameters required by the function, also point it out.
-You should only return the tool call in tools call sections.
-If you decide to invoke any of the function(s), you MUST put it in the format of:
-[[func_name1(params_name1=value1, params_name2=value2,…), func_name2(params)]]
-You should not include any other text in the response. Here is a list of functions in JSON format that you can invoke:
-{candidate tools}
-{other information}
-Figure 5: The system prompt for tool calling.
-“{candidate tools}” is replaced with actual tool descriptions.
-“{other information}” is replaced with some other information in the case, e.g., date.
-If no other information, it is null.
-Please refine your answer. Directly output the refined answer, or the original answer if you think it is already perfect.
-Figure 6: The refine prompt for tool calling.
-[System]
-You are an expert in composing functions.
-You are given a question and a set of possible functions.
-Based on the question, you will need to make one or more function/tool calls to achieve the purpose.
-If none of the function can be used, point it out.
-If the given question lacks the parameters required by the function, also point it out.
-You should only return the tool call in tools call sections.
-If you decide to invoke any of the function(s), you MUST put it in the format of:
-[[func_name1(params_name1=value1, params_name2=value2,…), func_name2(params)]]
-You should not include any other text in the response. Here is a list of functions in JSON format that you can invoke:
-    "name": "bookFlight",
-    "description": "Book a flight for a specified destination",
-    "parameters": {
-        "type": "dict",
-        "properties": {
-            "origin": {
-                "type": "string",
-                "description": "The departure airport or city"
-            "destination": {
-                "type": "string",
-                "description": "The destination airport or city"
-            "departure_date": {
-                "type": "string",
-                "description": "The date of departure (in YYYY-$M$-DD)"
-        "required": [
-            "origin",
-            "destination",
-            "departure_date"
-The current time is 2024-10-02 18:18:11.
-[User]
-I need to book a flight from New York to London for a business trip tomorrow.
-[Assistant]
-[[bookFlight(origin="New York", destination="London", departure_date="2024-10-03")]]
-[User]
-Please refine your answer. Directly output the refined answer, or the original answer if you think it is already perfect.
-[Assistant]
-[[bookFlight(origin="New York", destination="London", departure_date="2024-10-03")]]
-Figure 7: One example for refine data.
-In this example, the assistant does not change the tool invocation as it is correct already.
-Through this kind of data, the model is able to learn when to stop iteration during inference.